### v3

# Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps

Abstract.
In recent years, spatial computing Virtual Reality (VR) has emerged as a transformative technology, offering users immersive and interactive experiences across diversified virtual environments. Users can interact with VR apps through interactable GUI elements (IGEs) on the stereoscopic three-dimensional (3D) graphical user interface (GUI). The accurate recognition of these IGEs is instrumental, serving as the foundation of many software engineering tasks, including automated testing and effective GUI search. The most recent IGE detection approaches for 2D mobile apps typically train a supervised object detection model based on a large-scale manually-labeled GUI dataset, usually with a pre-defined set of clickable GUI element categories like buttons and spinners. Such approaches can hardly be applied to IGE detection in VR apps, due to a multitude of challenges including complexities posed by open-vocabulary and heterogeneous IGE categories, intricacies of context-sensitive interactability, and the necessities of precise spatial perception and visual-semantic alignment for accurate IGE detection results. Thus, it is necessary to embark on the IGE research tailored to VR apps.
In this paper, we propose the first zero-shot cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps, named Orienter. By imitating human behaviors, Orienter observes and understands the semantic contexts of VR app scenes first, before performing the detection. The detection process is iterated within a feedback-directed validation and reflection loop. Specifically, Orienter contains three components, including (1) Semantic context comprehension for capturing the apps’ GUI context, (2) Reflection-directed IGE candidate detection for identifying and localizing valid GUI elements based on multi-perspective description guided IGE detection, as well as feedback-directed reflection, and (3) Context-sensitive interactability classification which integrates semantic contexts for interactability prediction. To evaluate our approach and facilitate follow-up research, we spend more than three months constructing the first benchmark dataset which contains 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories. Extensive experiments on the dataset demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches (i.e., GPT-4o, YOLO v8, CenterNet2, Faster R-CNN, UIED, and Xianyu), surpassing their F1 Score by up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) in distinguishing the interactibility and semantics of the IGEs, respectively. Orienter is beneficial for boosting the performance of automatic testing by isolating the interactable action space from the whole space, regardless of the testing strategies employed. Experiments demonstrate that Orienter-guided testing covers 103.1% more IGEs with 125.7% more effective interactions than testing without action space isolation.
Virtual Reality, Spatial Intelligence, GUI Agent, Large Multimodal Model
ccs: Software and its engineering Software development techniquesccs: Computing methodologies Computer visionccs: Human-centered computing Virtual reality
1. Introduction
In recent years, Virtual Reality (VR) has emerged as a transformative technology, offering users immersive experiences across various virtual or virtual-real environments. This technological advancement has catalyzed the development of a myriad of VR apps, now numbering in tens of thousands (Li et al., 2023). These apps, which span a diverse array of domains, including skill training (web, 2021b), entertainment (web, 2021c, a), medical procedures (Khor et al., 2016), and military training (Bhagat et al., 2016), have attracted over 171 million users (Statista, 2022). This exponential growth underscores a critical need for robust software processes, including development, testing, and maintenance, particularly in high-reliability contexts such as healthcare and military training.
In VR apps, users experience multimodal perceptions through various devices and interact via body movements and gestures. Among the multimodal perceptions, visual perception obtained via an app’s Graphical User Interface (GUI) offers the wealthiest information. GUIs in VR apps are often composed of three-dimensional (3D) GUI elements (e.g., images, text, widgets, etc.) or real-life objects and users interact with VR apps through Interactable GUI Elements (short as IGEs).
Literature has shown that the accurate recognition of IGEs is a cornerstone for many software engineering tasks including automated testing (Xie et al., 2020; Ye et al., 2021; Chen et al., 2020) and effective GUI search (Reiss et al., 2018; Deka et al., 2017). For example, Ye et al. (Ye et al., 2021) report that 77% software experts believe that precise IGE detection can boost software testing efficiency by at least 50%. White et al. (White et al., 2019) demonstrate that IGE detection improves branch coverage by 42.5% compared to random testing.
Deep learning (DL) based object detection approaches like Faster R-CNN (Ren et al., 2017), CenterNet (Zhou et al., 2021) and YOLO (Jocher et al., 2023), have demonstrated promising performance. With such advancements, recent work has taken steps to explore DL-based IGE detection approaches in mobile apps and desktop apps (White et al., 2019; Chen et al., 2020; Xie et al., 2020; Ye et al., 2021). For example, Wu et al. (Wu et al., 2023) explore advanced DL-based approaches, like CenterNet2, YOLOv3, and YOLOv5, for IGE detection on mobile apps. The approaches can be typically divided into three steps: (1) manually summarize a finite set of IGE categories such as buttons, spinners, switches, checkboxes, etc., (2) manually label a large dataset, and then (3) train an object detection model on the dataset. However, such training-based approaches heavily rely on large annotated datasets so that they can hardly be directly applied to IGE detection in VR apps. As shown in Figure 1(a) and Figure 1(b), annotating large-scale VR IGE datasets is challenging due to the unique interaction mechanism of VR apps. The data annotation process demands extensive efforts, compounded by VR’s complex hardware usage and exhaustive interactions.
Recent advancements of pretrained large multimodal models (LMMs) (OpenAI, 2023; Gemini Team, 2023) have shown their remarkable abilities in downstream tasks including image/natural language comprehension, question answering and logical reasoning, without large training sets on specific problems. Such capabilities provide us with new opportunities to resolve the aforementioned limitation of lacking datasets. However, our preliminary experiments (detailed in  section 2.2) reveal that LMMs suffer from severe spatial semantic hallucinations, tending to generate contextually coherent but factually incorrect or unrealistic IGEs with inaccurate locations and amounts as responses. These are mainly stemmed from the following challenges of VR IGE detection problem:
Challenge #1: Open-vocabulary and heterogeneous GUI element categories. LMM or DL approaches perform well on detecting a finite pre-defined set of IGE categories that frequently appear in their training sets and adhere to standard visual appearance patterns (e.g., buttons, sliders, and checkboxes) as well as unified interaction mechanisms (e.g., tapping and long-tapping). However, VR apps with diverse scenarios contain infinite open-vocabulary categories of IGEs that own different visual appearances and interaction mechanisms (heterogeneity), as shown in Figure 1(a). In the stereoscopic 3D VR scene, from time to time, IGEs occlude each other or are observed from non-front perspectives (e.g., Figure 3(a)), making the visual appearances of even the same IGE overly diverse and heterogeneous. Hence, it is hard for LMM or DL approaches to identify such open-vocabulary and heterogeneous VR GUI element categories.
Challenge #2: Context-sensitive interactability. The interactability of GUI elements in VR apps is highly dependent on the semantic context of the VR app content and scenarios. As shown in Figure 1(c) and 1(d), respectively, an object like a tree might be interactable in one app (e.g., a tree-planting game where it needs to be picked up and planted) but non-interactable in another (e.g., as a mere background element in a fishing game). The IGE categories that LMM or DL approaches perform well when GUI elements with similar appearances behave mostly the same in distinct contexts, like buttons and checkboxes. We call them context-independent GUI elements. However, the interactability of GUI elements in VR apps is highly dependent on the app-specific and scenario-specific context, which means they cannot be uniformly determined by appearances but require context-sensitive reasoning instead.
Challenge #3: Accurate IGE detection results require precise spatial perception and visual-semantic alignment. While capable of processing visual data and generating textual descriptions, LMM struggles with accurate spatial perception. Their ability to outline and label objects is limited, often resulting in imprecise localization, misaligned bounding boxes, etc. This spatial deficiency hinders their applicability in VR environments, where precise IGE localization is necessary.
(a) IGE categories in a VR lab training app (web, 2024d) with diverse visual appearances and interaction mechanisms
(b) Diversified IGE categories in Figure 1(a) with annotations, dataset annotation requires massive effort
(c) An interactable tree in a VR gardening (tree planting) game (web, 2024e)
(d) Several non-interactable trees in a VR fishing game (web, 2024g)
Figure 1. Examples to demonstrate the challenges of IGE detection for VR apps (web, 2024e, g, d)
To tackle the challenges, we propose a cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps (short as Orienter). Instead of directly identifying the IGEs in VR apps, Orienter is designed to imitate human behavior by comprehending the semantic context first before conducting the detection; and the detection process is iterated within a feedback-directed validation and reflection loop (i.e., two levels of “looking before leaping” in the title). Specifically, Orienter mainly contains three major modules. ❶ Semantic context comprehension: For identifying heterogeneous IGE categories in an unsupervised and context-sensitive way, we make Orienter to understand and combine the global context (app content overview, including genres, storylines, interaction mechanisms, etc.) and local context (the current VR scenario within the user’s field of view) of the VR scenes under analysis through in-context reasoning of LMM. This phase is the foundation of all the subsequent steps. ❷ Reflection-directed IGE candidate detection: For identifying and localizing valid IGEs, Orienter first mines multi-perspective characteristics of GUI elements leveraging LMM, and then uses these characteristic descriptions as guidance to detect corresponding elements. Based on the detection results, Orienter further performs feedback-directed reflection to validate and refine IGE candidates using a feedback loop. ❸ Context-sensitive interactabilitiy classification: To predict the interactibility of detected IGEs, Orienter incorporates the semantic context from the first module and performs chain-of-thought in-context classification.
For facilitating the evaluation, we spend more than three months constructing the first benchmark dataset for IGE detection in VR apps. The benchmark dataset consists of 1,552 images from 100 industrial-setting apps on Steam (web, 2023c), with 4,470 interactable annotations across 766 semantics categories. We evaluate Orienter in terms of predictions on interactability and semantics. Orienter outperforms all baselines (GPT-4o (web, 2024b), YOLO v8 (Jocher et al., 2023), CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), UIED (Xie et al., 2020), Xianyu (Chen et al., 2020)) in almost all metrics, including precision, recall, and F1 score. Specifically, Orienter achieves improvements of up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) on F1 scores in distinguishing the interactability and semantics of IGEs, respectively, demonstrating Orienter’s strong effectiveness, efficiency, and reliability. To illustrate the usefulness of Orienter on downstream software engineering (SE) tasks, we further evaluate how effective it boosts the performance of automated testing. Orienter’s IGE detection results can be utilized to isolate the interactable action space from the whole input space, regardless of the testing strategies employed. Experiment results reveal that Orienter-guided testing covers 103.1% more IGEs with 125.7% more effective interactions than testing without action space isolation.
In summary, we make the following contributions in this paper:
  • To the best of our knowledge, we are the first to formulate the interactable GUI element (IGE) detection problem for stereoscopic 3D GUI. We leverage the power of large multimodal models in understanding and analyzing GUIs of VR software, and propose a zero-shot context-sensitive GUI element detection framework, Orienter. Orienter novelly undergoes semantic context comprehension, reflection-directed IGE candidate detection, and context-sensitive interactability classification to tackle the challenges of VR IGE detection.
  • Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework. The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, surpassing their F1 Score by up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) in distinguishing the interactibility and semantics of the IGEs, respectively. Experiments also illustrate that Orienter is beneficial for boosting GUI testing by covering 103.1% more IGEs with 125.7% more effective interactions than random testing.
  • We construct the first dataset for VR IGE detection, including 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories.
2. Preliminaries
2.1. Background of Virtual Reality (VR)
2.1.1. VR Devices
VR technologies provide immersive experiences that blend virtual elements with the physical world. These experiences are accessible via diverse devices including PCs and standalone systems, creating interactive environments that integrate real and virtual elements. Key VR devices include: (a) Head-mounted displays (HMDs), essential for VR experiences, offering immersive visuals and audio. (b) Two handheld controllers and gesture recognition systems enable users to interact with the virtual world via physical actions and gestures. (c) Tracking systems, comprising sensors and cameras, track user movement and orientation, crucial for a responsive VR experience.
2.1.2. VR Interaction
Interaction within VR environments can occur through various mechanisms: (a) Pressing buttons on controllers to manipulate GUI elements. (b) Touching or moving devices towards GUI elements, sometimes using extended virtual tools for remote interaction. (c) Gazing at GUI elements or employing eye-tracking technologies in HMDs to select or move elements. (d) Gestures and Movements to interact with the virtual environment through physical actions. These interaction modes highlight the evolving nature of human-computer interaction within VR, offering increasingly diverse and immersive experiences.
2.2. Motivating Examples
In this section, we present our motivational study to further demonstrate the technical challenges listed in section 1.
Two authors randomly sample 30 VR apps from the popular Steam VR app store (web, 2023c), select screenshots from them, and use the prompt template in Figure 2 to instruct the widely-used LMM GPT-4o (web, 2024b) to identify and locate IGEs in the images.
Figure 2. Propmt template for motivating examples.
As shown in figure 3(a), the mailbox is wrongly recognized as “Missile-like Object”. It is because the mailbox is viewed from an unusual angle, resulting in a special shape in the view and making it harder for the LMMs to identify semantics. This misunderstanding of GUI elements reveals the LMM’s inability to face the heterogeneous GUI elements in VR apps.
(a) W.r.t heterogeneity
(b) W.r.t. spatial perception
Figure 3. LMM gives deviated IGE detection results on VR scenes (web, 2024f, c) because of aforementioned challenges
For another example, figure 3(b), the bounding boxes in the image either cover only part of the object, locate objects mismatched with the label, or even mark the background. Although the LMM identifies most objects in the image, it fails to locate them well, revealing its insufficient spatial perception.
To summarize, the LMM alone cannot effectively detect IGEs in highly complex and varied VR apps. To boost the LMM on IGE detection in VR apps, it is necessary to provide the LMM with more context information and integrate the model with different techniques. Following this idea, we propose Orienter to imitate human behavior to enhance the LMM on detecting IGEs in VR apps.
3. The Orienter Approach
3.1. Problem Formulation of IGE Detection
Before jumping into the details of our approach, we first introduce how we formulate the IGE detection problem. The basic IGE detection unit is an individual VR app scenario, which lies in the user’s field of view, at a specific time. Unlike mobile application GUIs or web app GUIs, which are entirely 2D, HMD-based VR apps typically render two 2D GUIs for both eyes of users, creating an illusion of depth and making users feel stereoscopic 3D (S3D) sense (Penn and Hout, 2018; Elmqaddem, 2019). These scenes, presented to each eye, are projections of the 3D virtual world onto two 2D ”picture planes” (user’s eyes), mathematically represented by the function: $P_{3D\rightarrow 2D}:UI_{3D}\rightarrow(UI_{left},UI_{right}),$ where $UI_{3D}$ represents the 3D scene in the VR environment, and $UI_{left}$ and $UI_{right}$ represent the corresponding 2D projections rendered for the left and right eyes, respectively. Each one-eye scene can be recovered using the other eye’s scene (Li et al., 2024). Therefore, the S3D IGEs can be detected by analyzing the 2D projection from a single eye. This allows us to simplify the VR IGE detection problem to detecting IGEs in the 2D scene of any individual eye.
To facilitate downstream SE tasks such as boosting automated GUI testing, each IGE is identified not only by its interactability status but also by its location and associated semantic labels, enabling a more fine-grained understanding of the IGEs. Let $GE$ represent the set of GUI elements in the VR scene under analysis, i.e., $UI$. Given that the 3D scene $UI_{3D}$ is projected into 2D for each eye, the right-eye scene can be denoted as: $P_{3D\rightarrow 2D}(UI_{right}):UI_{3D}\rightarrow UI_{right}.$ The detection function $D:GE_{2D}\rightarrow\{Non-interactable,Interactable\}$ maps each GUI element to a binary interactability determination. The function $D$ operates on the 2D projections of the 3D GUI elements, denoted as:
$D(P_{3D\rightarrow 2D}(UI_{right})):C\times P_{3D\rightarrow 2D}(GE_{3D})\rightarrow\{Non-interactable,Interactable\}\times B\times S.$
Here, $C$ denotes the set of semantic contexts that facilitate context-sensitive analysis. We use a bounding box $B$ to specify the spatial location of each IGE, which is the smallest unrotated rectangle containing that IGE and can be described by its upper-left corner coordinates, width, and height. Let $S$ denote the set of semantic labels associated with each GUI element, where the labels and their granularity depend on the specific VR scene. As presented above, the problem can be simplified to detection on the 2D scene of any individual eye. If we use the right-eye scene for demonstration, the IGE detection problem can be formulated as:
$D(P_{3D\rightarrow 2D}(UI_{right})):C\times GE_{2D}\rightarrow\{Non-interactable,Interactable\}\times B\times S.$
Figure 4. Overview of Orienter. The complete prompt templates and example answers can be found on our website. The rounded rectangles in the figure represent modules and submodules. The arrows represent information flow. The light yellow boxes represent diverse forms of data.
3.2. Overview of Orienter
Figure 4 illustrates our cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps, short as Orienter. To address the three challenges (section 1), Orienter employs a feedback-driven approach that mirrors human behavior, prioritizing semantic comprehension before initiating detection; Orienter also enables iterative refinement and validation through a feedback loop. Such designs embody the principle of “looking before leaping”.
Orienter consists of three primary components: (1) Semantic Context Comprehension: To enable unsupervised, context-sensitive IGE detection, Orienter first synthesizes both the global context (e.g., app genre, storyline, interaction mechanisms) and the local context (i.e., the user’s current VR field of view). Leveraging the LMM, Orienter performs in-context reasoning to establish a comprehensive understanding of the VR scene, forming the basis for subsequent detection. (2) Reflection-Guided IGE Candidate Detection: Orienter then mines multi-perspective characteristics of GUI elements, and uses these characteristic descriptions for detecting and locating potential IGEs. The framework iteratively refines these candidates through a feedback loop, performing reflection-driven validation to reduce hallucination. (3) Context-Sensitive Interactability Classification: Finally, Orienter predicts the interactability of the detected IGE candidates by incorporating the semantic context from the first stage. Using chain-of-thought reasoning within the LMM, Orienter classifies interactability without requiring labeled training data, leveraging a knowledge-transfer paradigm from pretrained models in other domains.
3.3. Module I: Semantic Context Comprehension
The interactability of GUI elements in VR apps is highly dependent on the semantic context of the VR app content and scenarios. To tackle this challenge, Orienter understands and analyzes the target VR app’s context beforehand to enable context-sensitive detection. The semantic context is analyzed comprehensively, from (1) both global (overall app context) and local (current VR scenario) granularities, and (2) both natural language (texts) to vision (images) modalities.
3.3.1. Global Context Extraction.
The global context captures the overall semantic information of the VR app under analysis. To find out which global context attributes have influences on IGE interactability, two authors randomly sample 30 real-world VR applications from SteamVR and perform manual inspection following the open coding procedure (Creswell and Poth, 2016). At last, we reach a consensus and locate the following key attributes: app names, genres, content themes, VR device supports, ways of gameplay, possible interaction mechanisms, and language information. Such context can be extracted from the detailed information page on official VR app stores, but effective extraction can be troublesome due to the free structure natural language format of app details. Hence, Orienter leverages LMM to perform the extraction after the app details are crawled automatically. Prompt PI.1 in Figure 2 shows the prompt template.
Figure 5. Prompt templates for Module I [^2] Notes for all prompt templates: Red-colored properties are textual inputs from other modules, pink-colored are image inputs, orange-colored are outputs, and blue-colored are automatically generated information such as CoT demonstrations. We analyze 30 randomly selected SteamVR apps and write CoT reasoning steps and results as demonstrations in the template for LMM. Orienter randomly selects three of them as demonstration examples during runtime.
3.3.2. Local Context Perception.
Local context represents the semantic context within the current VR scenario, i.e., all VR content within the user’s current field of view. Under the extracted global context of app content from section 3.3.1, Orienter further guides LMM to perform in-context visual question answering, to digest and summarize the rendered VR scene as local context. The guidance of global context ensures that the interpretation of the VR scene is informed by the narrative and purpose of the VR app under analysis. Specifically, the LMM processes the composite input (screenshot and global context) and outputs a summary of all GUI elements and the background within the scene. Prompt PI.2 in Figure 2 shows the prompt template.
3.4. Module II: Reflection-Directed IGE Candidate Detection
This module aims at recognizing and localizing IGE candidates within the VR scene, based on the captured semantic contexts from Module I. LMM can reveal more semantic hints than traditional methods, making it a better backbone for our approach. However, LMM tends to generate semantically-incorrect IGEs with inaccurate locations. This is because VR IGE detection requires precise spatial perception and visual-semantic alignment capabilities, and VR IGE categories are open-vocabulary and heterogeneous (as presented in the challenges and motivational examples). Orienter conquer these challenges through a reflection-directed IGE candidate detection loop:
i). Since LMM has difficulties locating IGE candidates and produces semantic hallucinations, Orienter performs IGE candidate detection (section 3.4.2) upon LMM’s results, based on a language-model-aligned visual foundation model (VFM). This VFM detection module (short as VFMD) verifies IGE candidates’ existence, reduces hallucinations of incorrect candidates, and locates the exact bounding boxes of existing candidates for further analysis.
ii). Then the VFMD module and LMM work as a chain to conduct IGE detection. Although VFMD can help reduce hallucinations from LMM, its own analysis process is still error-prone. To reduce errors in the two components simultaneously, Orienter performs iterative validation and refinement alongside the reflection loop (section 3.4.3). During validation, Orienter let another LMM agent, as an advisor (section 3.4.3), compare and rethink all the discrepancies and consistencies between the results from the LMM detector (section 3.4.1) and VFMD module. If any concerns exist for some IGE candidates, the detection chain will run again with the validation comments from the LMM advisor. The whole process ends when the LMM advisor claims confidence for all detection results.
iii). Inspired by the human cognitive process, when trying to find and locate IGE candidates, using a semantic label only will make the detection process unrobust, e.g., may find the wrong element or find the wrong locations. Adding more descriptions about the target element (like color, size, shape, relevant locations, etc.) makes the detection more effective and robust. It is similar for To boost the effectiveness and robustness of i) and ii), Orienter extends a direct LMM detector to a multi-perspective characteristics miner (section 3.4.1).
Overall, Orienter first mines multi-perspective characteristics of GUI elements leveraging LMM, and then uses these characteristic descriptions as guidance to detect corresponding elements. Based on the detection results, Orienter further performs feedback-directed reflection to validate and refine IGE candidates using a feedback loop.
3.4.1. Multi-Perspective Characteristics Mining
Figure 6. Example input for Module II.①, from the Baseball Kings VR app (web, 2024a)
In this module, Orienter firstly identifies all IGE candidates in the VR scene, and then mines their diverse characteristics from different perspectives (e.g., size, color, shape, relevant location with other objects, etc.). Then these characteristics will form a characteristics description, which can uniquely identify the corresponding IGE candidate in the present figure in the described detection module. Figure 7 shows the prompt templates and examples.
Figure 7. Prompt templates and examples for Module II.①2
II.①(a): IGE Candidate Recognition (Prompt PII.1). Orienter engages in the recognition of IGE candidates within the VR scene. With the semantic context from Module I (section 3.3), Orienter analyzes the visual semantic cues presented in the environment, finding out all IGE candidates.
II.①(b): Characteristics Dimension Recognition (Prompt PII.2). Following the identification of IGE candidates, Orienter embarks on recognizing the characteristics dimensions of them, which makes them distinguishable from other GUI elements. This involves dissecting the unique characteristics attributes and properties related to the visual appearance (e.g., color, size, shape, etc.), functionalities, interaction mechanisms, and relative locations of each GUI element.
II.①(c) Characteristics Question Formulation (Prompt PII.3). Merging results from the previous two steps, Orienter formulates multiple characteristic questions (one question for each dimension of characteristics) for each IGE candidate. These questions are tailored to probe the specific appearance, functionalities, relative locations, and interactability dimensions of GUI elements.
II.①(d) In-Context Characteristics Reasoning (Prompt PII.4). Then Orienter conducts characteristics reasoning, querying the LMM to analyze the VR scene and find out the answers for formulated characteristic questions. This step integrates the local and global contexts previously established, allowing the framework to interpret the responses within the broader narrative and functional scope of the VR application. Through this in-context reasoning, Orienter synthesizes a comprehensive understanding of each IGE candidate’s characteristics. Orienter then generates characteristics descriptions (CDs), which describe the GUI elements based on these characteristics. Figure 7 shows an example of a characteristics description, “A white and black small spherical baseball with grey shadows on the bottom half and glowing trail across the center of the scene”.
3.4.2. Described IGE Candidate Detection
Figure 8. The framework of described IGE candidate detection module (VFMD)
Figure 9. Example: described IGE candidate detection under the CD “donut with colored granules on the surface”, the donut in the middle is detected.
Orienter detects and localizes IGEs using the descriptions of the crafted characteristics generated in the previous step, inspired by Shen’s work (Shen et al., 2024). Orienter employs a language-model-aligned visual foundation model (VFMD). Figure 9 shows the framework structrure. VFMD consists of a vision backbone for extracting visual features from VR scenes and a language model for generating text embeddings from CDs. Text embedding s are aggregated into sentence-level representations and are fused with visual embeddings using a cross-modality transformer encoder. Object queries conditioned on the text are processed by a transformer decoder, producing object embeddings. A final visual-language alignment module predicts the correct pairings of IGE candidate locations and CDs. The backbone language model, trained on a vast corpus of data, can generalize to the diverse and unlimited categories of GUI elements present in VR scenarios. The localization process cross-references (or aligns with) the textual CDs with the visual input from the VR scene. By doing so, VFMD successfully locates IGE candidates. Here is an example to demonstrate the output of this module, as shown in Figure 9.
3.4.3. Feedback-Directed Reflection
Figure 10. Prompt templates and examples for Module II.③2
Following previous detection attempts, Orienter validates and refines the identified IGE candidates through a feedback-directed reflection loop. Orienter tries to compare and rethink both the discrepancies and consistencies between the detection results of VFMD module (Module II.②) and LMM characteristics miner (Module II.①). Specifically, Orienter distinguishes successfully and unsuccessfully detected CDs, and conducts mirroring-based regional verification and region-wise reflection respectively. Another LMM agent, as an advisor, analyzes the verification and reflection results and decides whether there exist any concerns for any IGE candidates (Prompt PII.7). If yes, go back to Multi-Perspective Characteristics Mining module to perform mining again, with the verification results. Orienter iterates through these steps until an LMM advisor decides it reaches sufficient confidence level for all detection results. Upon reaching a sufficient satisfactory level of confidence in the accuracy of all detected IGE candidates, Orienter progresses to the final phase of Context-Sensitive Interactability Classification. Figure 10 shows prompt templates and examples.
Region-Wise Reflection on Unsuccessfully Detected CDs (Prompt PII.6). We provide the undetected characteristics descriptions and the visualized IGE detection results to LMM to validate whether the non-detection is a result of model hallucination or a possible missing detection of the described IGE candidate detection module. If LMM decides that it is likely to be a detection miss, Orienter will revisit the characteristics mining phase to refine the characteristics descriptions, ensuring a more directed and precise subsequent detection attempt.
Mirroring-Based Regional Verification for Successfully Detected CDs (Prompt PII.5). Conversely, for successfully detected IGEs, we design a method of regional zooming and mirroring. This involves cropping the detected GUI element within its bounding box into a new figure, and Orienter lets LMM conduct a comparative analysis against the original VR scene. Such an approach facilitates a granular verification of the detection’s accuracy.
3.5. Module III: Context-Sensitive Interactability Classification
Figure 11. Prompt templates for Module III2
After establishing the presence and location of GUI elements in the VR environment through the first two modules, Orienter proceeds to analyze their interactivity. This module determines which GUI elements in the VR scene can be interacted with. Such information is useful for many downstream SE tasks, including boosting automated testing performance (detailed in section 6.3).
Figure 11 shows the prompt template PIII. To classify interactability, Orienter guides LMM towards a chain-of-thought in-context reasoning process, within the semantic context captured in Module I. This process mimics a VR player’s thought pattern in distinguishing between GUI elements that are mere visual elements and those that afford interaction. For instance, a tree in a tree-planting game (Figure 1(c)) is interactable, serving as a GUI element to be manipulated by the player, whereas the same tree would be non-interactable background scenery in a fishing game (Figure 1(d)). Orienter also provides demonstration examples in the prompt, illustrating the reasoning required for differentiating between interactable and non-interactable GUI elements.
4. Dataset Construction
To the best of our knowledge, there is no existing GUI dataset for IGE detection in VR apps. Therefore, we build a dataset to verify the usefulness and effectiveness of Orienter.
We recruit a team of 13 annotators, spending more than three months in the collection and annotation of GUI images from various VR apps. These annotators have a minimum of two years of computer science or electrical engineering background, and mostly with experience in video games. The annotators first interact with GUI elements in various VR apps while their views are recorded. They then select GUI images from these recordings and annotate IGEs. To enhance the quality of the dataset, we train the annotators in the use of VR equipment and the labeling of IGEs, and we ask them to follow several guidelines during data collection and annotation: (1) attempt to identify all IGEs using every possible interaction method in each VR scene, and (2) categorize the semantics of IGEs with appropriate granularity based on their context.
4.1. Collection of VR Apps
We first collect VR apps from the Steam app store (web, 2023c), a comprehensive repository with a wide variety of VR content. This collection yielded a total of 4,610 VR-Only apps, which includes software (official type label on Steam referring to non-games) and games. To obtain a statistically significant sample size with a 95% confidence level and a 10% margin of error, we randomly sample 102 apps for analysis to cover a broader spectrum of categories, contexts, and interaction paradigms, covering 245 community-generated genres on Steam, demonstrating their diversity and representativeness.
4.2. Collection of GUI Images from VR Apps
In this step, the annotators engage with the selected VR apps, interacting with all GUI components within the VR scene in every possible way, as described in section 2.1.2, while recording their stereo view. The annotators then select and save GUI images that encompass all VR scenes and GUI elements they explored from the recordings. We crop the right-eye images from the stereo-view recordings, as only one side of the view is needed, as explained in section 3.1.
(a) Cooking (web, 2023d)
(b) Big-eat-small (web, 2023b)
Figure 12. Examples to demonstrate varying granularity to best categorize semantics of IGEs.
4.3. Annotation of IGEs in GUI Images
In this step, the annotators identify IGEs in GUI images and label their locations with bounding boxes as well as their semantics within the context. As discussed in section 1, the interactability and semantics of GUI elements heavily depend on the diverse contexts of VR apps, making it challenging to categorize all GUI elements with a finite set of predefined categories.
To address this challenge, we ask the annotators to create categories. Specifically, they are instructed to classify IGEs with appropriate granularity based on their experience, to better describe the semantics of IGEs in concrete contexts. For example, in the VR game, VR The Diner Duo (web, 2023d), fish is only one kind of ingredient to make burgers, as shown in 12(a). Different individuals of fish are semantically the same in this game. Therefore, a coarse-grained fish category is adequate to describe it. However, in the game Munch VR (web, 2023b), players need to control a fish to eat other smaller fish while avoiding hunted by larger fish, as shown in 12(b). In this scenario, fish individuals diverge semantically according to size and appearance, necessitating the introduction of more fine-grained categories such as large fish and small fish to capture their semantic distinctions.
4.4. Dataset Statistics
After data cleaning, we finally construct a dataset consisting of 1,552 images with the size of 960 * 540 from 100 apps, covering 245 community-generated genres on Steam, with 4,470 interactable annotations across 766 semantics categories, highlighting the diversity of our dataset. Annotations are in widely used COCO (Lin et al., 2014) format, which uses bounding boxes to mark the locations of IGEs.
5. Experiment Design
5.1. Research Questions
In this study, our experiment is designed to answer the following research questions:
  • RQ1 (Performance in industrial-setting): How effective is our proposed framework, Orienter, in IGE detection on industrial-setting VR apps?
    • RQ1-1 (Interactability): How effective is Orienter in terms of analyzing interactability?
    • RQ1-2 (Semantics): How effective is Orienter in terms of inferring semantics?
    • RQ1-3 (Context-sensitive interactability): How does Orienter perform in terms of analyzing context-sensitive interactability?
  • RQ2 (Ablation Study): How does each component of Orienter contribute to its performance?
  • RQ3 (Usefulness): How effectively can Orienter boost automated testing on VR apps?
5.2. Baselines
Figure 13. Direct prompting
GPT-4o (web, 2024b) & Gemini 1.5 Pro (Reid et al., 2024) (direct prompting): Simply prompt the model with the image and ask it to locate all IGEs with bounding boxes and their semantics in the image to illustrate the end-to-end performance of LMM. Xianyu (Chen et al., 2020) is a UI-to-Code tool developed by Alibaba, leveraging old-fashioned computer vision techniques and OCR. UIED (Xie et al., 2020) combines old-fashioned methods with deep learning models to detect clickable GUI elements in complex GUI images. CenterNet2 (Zhou et al., 2021) is a two-stage object detector that uses class-agnostic one-stage detectors as the proposal network. It estimates object probabilities in the first stage and conditionally classify objects in the second stage. Faster R-CNN (Ren et al., 2017) is a two-stage anchor-based deep learning object detector. It contains a region proposal network to extract regions of interest. Objects within the RoIs are then classified with another neural network. YOLO v8 (Jocher et al., 2023) is a one-stage anchor-free detection model. It detects and classifies objects simultaneously and thus is faster than two-stage detectors.
5.3. Implementation Details and Experimental Setup
5.3.1. Implementation of Orienter
For the implementation of Orienter, we leverage the most representative and popular LMMs, i.e., gpt-4-vision-preview, gpt-4o-2024-08-06, Claude 3.5 Sonnet, Gemini 1.5 Pro. For described IGE detection, we leverage the pretrained model of APE-L_D (Shen et al., 2024), which is trained on ten datasets and demonstrates promising results in visual grounding. We apply post-processing to improve Orienter’s prediction quality by filtering abnormally large bounding boxes (over 90% of the image size), and applying Non-Maximum Suppression (NMS) which retains only the highest-confidence box among overlapping ones with IoU exceeds 0.7 to reduce duplicates.
5.3.2. Dataset Preparation
To comprehensively answer the RQs, we derive three variations from the original dataset. (1) Semantics dataset: The original dataset testing methods’ ability to identify IGEs’ semantics, answering RQ1-2. (2) Interactability dataset: All annotations are assigned the category ”interactable” for binary classification, testing models’ ability to differentiate IGEs, answering RQ1-1. (3) Context dataset: Contains 41 categories randomly sampled from the most common 100 categories, and extra annotations marking their corresponding non-interactable objects, testing methods’ understanding of context-sensitive IGEs in different contexts, answering RQ1-3.
We partitioned the images into training/validation/testing datasets in a 6:1:3 ratio using three distinct methods: App split, Genre split, and Context-sensitive split. The App split randomly allocates images based on their corresponding apps. The Genre split further considers the apps’ genres to assess the methods’ performance across different app types. The Context-sensitive split assigns images containing the 41 sampled categories in the context dataset to the test set while randomly distributing the rest into the training and validation sets.
5.3.3. Experimental Setup
We train baselines on the training dataset and evaluate all methods including Orienter on the test dataset. For Xianyu (Chen et al., 2020), we only use its component detection part to find IGEs. For UIED (Chen et al., 2020), we reimplement and train the CNN classifier on our dataset to adapt it to our problem, and replace the OCR component with PaddleOCR. For training on CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), and YOLO v8 (Jocher et al., 2023), we set the batch size to 32 and accordingly adjust the learning rate linearly, and keep other configurations the same as their original releases. We apply early stopping with the patience of 20 epochs to avoid overfitting. For LMM experiments, we run three times and regard the average results as final results.
We implement a semantic matching tool to support open vocabulary category matching. We consider two categories semantically match if the cosine similarity of their embedding vectors obtained from the embedding-3 model released by Zhipu AI (Zhipu AI, 2024) exceeds a preset threshold. The first two authors examine the similarity of some typical categories and set the threshold as 0.85.
For RQ1-1, RQ1-2, and RQ2, predictions are evaluated with our customized COCO API that adapts to our semantic matching tool. Note that during evaluation, instead of simply ignoring categories without ground truth annotations (usually due to splitting the dataset), as the official COCO API (web, 2023a) does, we further check if any predictions fall into those categories. Metrics of categories that meet this condition are set to 0 and included when averaging metrics across categories, providing more comprehensive results. Regarding RQ1-3, the baseline methods are trained on the Semantic dataset using Context-sensitive Split. Predictions on the test set are then evaluated on the Context dataset with the same split. The more IGEs and less non-interactable objects the method detects, the better this method’s performance on context-sensitive interactability understanding. Metrics are initially calculated for each category and then averaged to obtain the final results.
5.3.4. VR App Automated Testing Setup
To answer RQ3, we simulate a simplified test scenario: given a screenshot of the scene in the VR app under test, the testing agent attempts to interact with the GUI elements in the scene in a black-box setting. We compare the performance of the testing agent with and without the guidance of Orienter using the test set of Genre split as input.
We simplify the interaction events in the 3D virtual space to interaction points on UI images of VR scenes and ignore the type of interactions. The testing process is modeled as incrementally generating points on UI images over time, treating each point as an interaction made by the testing agent with an interval of one minute. We follow the recent works (Pan et al., 2020; Ghorbani et al., 2023) to set the testing duration to 60 minutes, which is longer than exploring a VR scene usually requires. We perform 5 testing runs and average the metrics to produce comprehensive results for each interaction strategy.
For the non-guided strategy, the testing agent randomly generates interaction points on the whole image. For the guided strategy, in each attempt, the testing agent generates interaction within the bounding boxes predicted by Orienter at a probability $p$, or ignores the prediction’s constraint to randomly explore the whole UI images at the probability $(1-p)$. The probability $p$ decreases gradually as the test progresses. Suppose the current moment is $t$ minutes and the total duration of the test is $T$ minutes, then $p=1-t/T$. This strategy allows the agent to generate effective yet diversified interaction events to cover more IGEs precisely.
5.4. Evaluation Metrics
For RQ1 and RQ2, we evaluate each method using Precision, Recall, F1-Score, and mAP. The calculation involves computing IoU.
IoU. Intersection over Union (IoU) is calculated as the ratio of two bounding boxes’ intersection area to their union area, measuring how well they match. A predefined threshold is usually set, and only pairs of bounding boxes with IoU above this threshold can match.
Precision, Recall, and F1-Score. A predicted bounding box matches the one with the highest IoU among the ground-truth bounding boxes that (1) lie in the same image, (2) have matching categories, and (3) have IoU exceeding the threshold. For RQ1-1, RQ1-2, and RQ2, we consider a predicted box matching a ground-truth box a True Positive (TP); otherwise, a False Positive (FP); ground-truth boxes that fail to match any predicted box are False Negatives (FN). For RQ1-2, a predicted box is considered a TP if it matches an interactable ground-truth box, and an FP if it matches a non-interactable one; non-interactable and interactable ground-truth boxes that fail to match any predicted box are considered a TN and FN, respectively. We calculate Precision as $TP/(TP+FP)$, Recall as $TP/(TP+FN)$ and F1-Score as $2\times Precision\times Recall/(Precision+Recall)$.
AP and mAP. We follow COCO API (web, 2023a) to calculate the Average Precision (AP) by averaging 101 Precision values in increments of 0.01 over a range of Recall values from 0 to 1. The mean AP (mAP) is obtained by averaging AP across categories.
We follow the previous works (Wang, 2022; Wang et al., 2023b) to use IGE Coverage and Effective Interaction Count in RQ3. Note that we can not calculate code coverage metrics in the black-box setting of RQ3.
IGE Coverage. The IGE coverage reflects the effectiveness of interaction strategies to find IGEs over time. An IGE is covered if at least one interaction point falls within the bounding box of that IGE. IGE Coverage is calculated as $n_{covered}/n_{all}$, where $n$ represents the number of IGEs.
Effective Interaction Count. The Effective Interaction Count reflects the efficiency of interaction strategies to make effective interactions by focusing on IGEs only. An interaction point is considered effective if it falls in any of IGEs’ bounding boxes.
6. Results and Analysis
6.1. RQ1: Performance in industrial-setting
6.1.1. RQ1-1: Performance in terms of interactability
As shown in Table 1, Orienter shows consistent performance in interactability detection across different IoU thresholds. Consider the generally best model Orienter with Gemini, at the 0.75 threshold, the mAP peaks at 35.03%, indicating a strong ability to detect IGEs accurately. However, as the IoU threshold increases, indicating a stricter criterion for object detection, there is a noticeable decrease in Precision, dropping to 5.64% at an IoU of 0.95. This trend suggests that while the Orienter is quite effective at a broader detection scope, its precision in highly specific object identification contexts is limited. Orienter demonstrates a balanced performance at lower IoU thresholds, with a peak F1 Score of 46.00% at an IoU of 0.75. This balance is crucial for practical applications, as it indicates a well-rounded capability in both correctly identifying IGEs and minimizing false positives.
For the comparative evaluation, Orienter generally demonstrates superior performance across IoU thresholds when compared with baseline models, especially on the genre split of the dataset, highlighting Orienter’s capability to handle various genres of apps. At an IoU threshold of 0.75, Orienter achieves a Precision of 49.45%, which is approximately 48.0% higher than the best baseline, Faster-RCNN (Ren et al., 2017), at 33.42%. This trend of superiority is more significant at higher IoU thresholds. At the stringent IoU of 0.95, Orienter maintains a Precision of 19.02%, outperforming the closest baseline, CenterNet2 (Zhou et al., 2021), by over 321.7%, a significant margin considering the complexity of VR environments. In terms of Recall and F1 Score, Orienter significantly outperforms other models by considerable margins. At an IoU of 0.75, Orienter’s Recall and F1 Score are 43% and 46.00%, respectively, surpassing CenterNet2 (Zhou et al., 2021)’s 30% and 29.32% by a notable margin of 43.3% and 56.9%, respectively. This indicates that Orienter is not only accurate but also reliable in identifying IGEs, a key requirement in industrial VR apps. Note that the direct prompt method on LMMs performs poorly. This is because although the LMM can identify the objects well, it can not effectively localize them, as discussed in section 2.2. This result again reveals the inability of LMM alone to address the challenge in IGE detection, highlighting Orienter’s enhancement for LMM on this task.
6.1.2. RQ1-2: Performance in terms of semantics
Orienter’s performance in semantics shows a similar but lower pattern compared to interactability. For Orienter with Gemini, the mAP starts at a lower rate of 22.12% at an IoU of 0.75 and follows a smooth decreasing trend with increasing IoU thresholds until 0.95, with a drop to 7.41%. This pattern reflects the challenges inherent in semantic interpretation, especially under stringent detection criteria.
The Precision and Recall metrics provide further insights. While the Recall maintains at 50.00%, the Precision starts at 23.72% at an IoU of 0.75 and gradually decreases. This suggests that Orienter maintains its ability to identify relevant semantic elements. However, Orienter faces challenges in high precision scenarios, particularly as the criteria become more stringent.
As for the comparative evaluation in terms of semantics, the results in Table 1 reveal Orienter’s remarkable capability in semantic understanding. At an IoU of 0.75, Orienter achieves a Precision of 23.72% with Gemini on the genre split, which is significantly higher than the best baseline CenterNet2 (Zhou et al., 2021)’s 0.62%, marking an increase of 3725.8%. Similarly, at an IoU of 0.95, Orienter’s F1 Score of 14.31% far surpasses the closest baseline, YOLO v8 (Jocher et al., 2023), which scores 0.40%. This represents an increase of up to 3477.5%, highlighting Orienter’s ability to understand complex semantic structures within VR environments. The direct prompting method continues to perform poorly in terms of semantics due to the aforementioned challenges that LMM alone can not effectively address. Note that the method directly prompting Gemini achieves a recall of 50% on the genre split at 0.75 IoU threshold. This result suggests that LMM can perform well in some special settings.
Our analysis demonstrates that Orienter significantly outperforms existing baseline models in both interactability and semantic understanding. The framework exhibits remarkable precision and recall rates, particularly in high IoU thresholds, which are critical for the nuanced and complex nature of industrial VR apps. This efficacy, especially in semantic understanding, underscores Orienter’s potential in revolutionizing IGE detection in VR environments.
6.1.3. RQ1-3: Performance in terms of context-sensitive interactability
As shown in table 2, Orienter shows a promising performance to distinguish IGEs’ interactability in different contexts. Orienter with Gemini performs better in terms of Precision, which achieves 67.91% at the 0.75 IoU threshold and maintains superior performance over 65% as the threshold increases, until at IoU threshold of 0.95, dropping to 53.35% which is still relatively high. Orienter with GPT-4o shows a better performance in terms of Recall and F1 score, achieving 24.00% and 32.51% at the 0.75 IoU threshold, demonstrating its better balance between Precision and Recall. Regarding the comparison with baselines, Orienter shows superior performance against baseline methods across different IoU thresholds. Consider the generally better model, Orienter with GPT-4o, at the lower IoU threshold of 0.75, Orienter achieves an F1 score of 32.51%, surpassing the best baseline CenterNet2 (Zhou et al., 2021) by 6274.5%. At the most stringent constraint of 0.95 IoU threshold, Orienter’s F1 score maintains over 10% while all metrics of baseline models are approximately 0, demonstrating Orienter’s significant advantage in understanding context-sensitive IGEs’ interactability against baseline models.
Table 1. Performance of Orienter, DL-based baselines are trained on our dataset*
Methods Interactability Semantics
App Split Genre Split App Split Genre Split
IoU 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95
Faster- RCNN mAP18.05 13.84 7.32 1.47 0.07 17.90 12.88 6.96 1.56 0.13 0.21 0.15 0.03 0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
P33.41 32.29 23.04 11.08 2.15 33.42 29.55 20.10 8.06 1.14 0.91 0.83 0.38 0.16 0.06 0.07 0.06 $\approx$0.0 $\approx$0.0 $\approx$0.0
R32.00 25.00 18.00 8.00 1.00 29.00 22.00 15.00 6.00 1.00 4.00 4.00 3.00 3.00 4.00 2.00 2.00 $\approx$0.0 $\approx$0.0 $\approx$0.0
F1% 32.69 28.18 20.21 9.29 1.36 31.05 25.22 17.18 6.88 1.06 1.48 1.37 0.68 0.30 0.11 0.13 0.12 $\approx$0.0 $\approx$0.0 $\approx$0.0
Center- Net2 mAP17.79 14.98 11.73 5.91 0.93 15.63 12.53 9.07 4.23 0.45 0.20 0.18 0.15 0.05 $\approx$0.0 0.22 0.19 0.05 0.02 $\approx$0.0
P32.53 28.52 26.85 18.26 4.83 28.67 25.81 27.71 16.46 4.51 0.68 0.65 0.57 0.27 0.05 0.62 0.59 0.27 0.19 0.04
R32.00 29.00 23.00 15.00 3.00 30.00 27.00 18.00 12.00 3.00 4.00 4.00 3.00 3.00 2.00 5.00 5.00 3.00 3.00 3.00
F1% 32.26 28.76 24.77 16.47 3.70 29.32 26.39 21.82 13.88 3.60 1.16 1.12 0.96 0.49 0.10 1.10 1.06 0.49 0.35 0.08
Yolo v8 mAP15.30 13.12 9.85 6.16 1.51 11.54 10.02 7.59 3.86 0.46 0.02 0.02 0.01 $\approx$0.0 $\approx$0.0 0.03 0.02 0.02 0.01 0.01
P26.58 25.38 20.63 16.65 7.31 23.92 21.38 24.22 14.55 2.68 0.21 0.20 0.19 0.05 0.01 0.29 0.28 0.27 0.25 0.21
R26.00 23.00 20.00 14.00 3.00 23.00 22.00 15.00 11.00 3.00 4.00 4.00 4.00 1.00 3.00 3.00 3.00 3.00 3.00 3.00
F1% 26.29 24.13 20.31 15.21 4.25 23.45 21.69 18.53 12.53 2.83 0.40 0.38 0.36 0.09 0.02 0.53 0.52 0.50 0.46 0.40
[…]
P0.21 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.43 0.43 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.25 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.23 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
[…]
F1% $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.47 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.46 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
gpt-4o- 2024-08-06 mAP0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
P0.78 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.75 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
[…]
F1% $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
Orienter (gpt-4-vision- preview) mAP25.53 24.32 23.03 19.64 4.61 27.07 25.22 23.87 19.80 4.54 14.93 14.66 13.96 12.17 3.67 15.96 15.57 14.62 12.87 3.75
P34.79 33.52 32.32 29.31 14.92 38.42 36.70 34.97 31.40 16.79 16.06 16.05 15.17 15.27 6.53 17.11 16.85 15.78 15.93 6.14
R40.00 39.00 38.00 34.00 12.00 40.00 38.00 37.00 32.00 11.00 50.00 50.00 50.00 33.00 20.00 50.00 50.00 50.00 33.00 20.00
F1% 37.21 36.05 34.93 31.48 13.30 39.20 37.34 35.95 31.70 13.29 24.31 24.30 23.28 20.88 9.84 25.50 25.21 23.98 21.48 9.39
Orienter (gpt-4o- 2024-08-06) mAP26.44 25.26 23.98 20.35 4.48 26.11 24.50 23.51 19.42 4.38 18.00 17.65 16.91 14.28 5.75 19.45 18.97 18.07 15.40 5.68
P34.63 34.89 35.43 31.64 18.57 37.91 36.68 36.71 33.24 12.91 19.32 18.97 18.28 15.37 6.87 20.35 19.81 19.01 16.19 7.36
R42.00 39.00 36.00 34.00 11.00 41.00 39.00 37.00 33.00 14.00 50.00 50.00 50.00 50.00 42.00 50.00 50.00 50.00 50.00 25.00
F1% 37.96 36.83 35.71 32.78 13.81 39.39 37.80 36.85 33.12 13.43 27.87 27.50 26.77 23.52 11.81 28.93 28.38 27.55 24.46 11.37
Orienter (Claude 3.5 Sonnet) mAP23.99 22.90 21.54 18.46 3.90 25.50 23.96 22.21 18.49 3.84 16.14 15.79 15.17 12.73 5.01 19.01 18.51 17.78 15.00 6.67
P31.73 37.52 36.79 33.56 16.30 37.97 40.15 39.22 35.86 14.51 16.63 16.28 15.83 13.35 7.64 19.68 19.11 18.53 15.75 9.87
R40.00 32.00 31.00 29.00 11.00 37.00 33.00 32.00 29.00 13.00 50.00 50.00 50.00 50.00 20.00 50.00 50.00 50.00 50.00 20.00
F1% 35.39 34.54 33.65 31.11 13.14 37.48 36.23 35.24 32.07 13.71 24.95 24.56 24.04 21.07 11.06 28.24 27.66 27.04 23.96 13.22
Orienter (Gemini 1.5 Pro) mAP32.71 31.27 30.51 26.33 5.88 35.03 32.59 30.82 25.08 5.64 20.69 20.35 20.19 17.77 7.16 22.12 21.44 21.17 18.45 7.41
P46.42 45.30 44.94 41.64 21.46 49.45 47.52 46.42 41.94 19.02 22.64 22.29 22.16 19.39 9.66 23.72 22.86 22.64 19.58 8.35
R42.00 41.00 40.00 37.00 14.00 43.00 42.00 41.00 37.00 15.00 50.00 50.00 50.00 50.00 25.00 50.00 50.00 50.00 50.00 50.00
F1% 44.10 43.04 42.33 39.18 16.95 46.00 44.59 43.54 39.32 16.77 31.17 30.84 30.71 27.94 13.94 32.17 31.38 31.16 28.14 14.31
    The highest metrics are colored with gray. Metrics below 0.01% are regarded as $\approx$ 0.0. All metrics of UIED and Xianyu are lower than 0.01% and omitted.
Table 2. Performance w.r.t. context-sensitive interactability, DL-based baselines are trained on our dataset*
Faster-RCNN CenterNet2 Yolo v8 gpt-4o-2024-08-06 Gemini 1.5 Pro
Orienter
(gpt-4-
vision-preview)
Orienter
(gpt-4o-
2024-08-06)
Orienter
(Claude 3.5
Sonnet)
Orienter
(Gemini 1.5 Pro)
IoU PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1%
0.75 2.44 0.14 0.27 4.27 0.27 0.51 $\approx$0.0 $\approx$0.0 $\approx$0.0 2.44 0.07 0.13 $\approx$0.0 $\approx$0.0 $\approx$0.0 44.72 6.66 10.77 67.70 24.00 32.51 62.74 18.54 26.07 67.91 22.96 31.40
0.80 2.44 0.09 0.18 4.88 0.27 0.52 $\approx$0.0 $\approx$0.0 $\approx$0.0 2.44 0.07 0.13 $\approx$0.0 $\approx$0.0 $\approx$0.0 44.57 6.32 10.41 67.99 23.36 31.82 62.93 17.94 25.36 67.91 22.20 30.63
0.85 $\approx$0.0 $\approx$0.0 $\approx$0.0 4.88 0.23 0.43 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 45.02 6.27 10.40 67.82 22.21 30.62 64.17 16.84 24.27 68.42 21.40 29.75
0.90 $\approx$0.0 $\approx$0.0 $\approx$0.0 4.88 0.18 0.34 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 45.24 5.37 9.26 68.93 19.35 27.75 63.96 14.84 21.84 70.01 18.23 26.78
0.95 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 26.83 1.62 3.03 47.56 8.99 13.59 47.07 7.47 11.37 53.35 7.01 11.76
Table 3. Contributions of different components*
Orienter w/o Context Comprehension w/o Reflection-Directed Loop w/o Interactability Classification
Interactability Semantic Interactability Semantic Interactability Semantic Interactability Semantic
IoU PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1%
0.75 47.14 43.00 44.97 21.36 50.00 29.93 38.27 39.00 38.63 18.14 50.00 26.62 38.51 36.00 37.21 18.85 50.00 27.38 31.40 29.00 30.15 16.78 50.00 25.13
0.80 45.95 42.00 43.88 20.82 50.00 29.40 35.80 39.00 37.33 17.71 50.00 26.15 36.18 36.00 36.09 18.34 50.00 26.83 29.63 29.00 29.31 16.47 50.00 24.78
0.85 45.37 41.00 43.08 20.64 50.00 29.22 34.90 38.00 36.38 17.13 50.00 25.51 36.36 34.00 35.14 17.57 50.00 26.01 29.91 27.00 28.38 18.90 33.00 24.04
0.90 41.76 37.00 39.23 17.69 50.00 26.14 33.00 33.00 33.00 17.13 33.00 22.55 32.87 31.00 31.91 14.51 50.00 22.49 26.77 25.00 25.86 16.46 33.00 21.97
0.95 19.25 16.00 17.47 8.18 50.00 14.06 13.80 14.00 13.90 6.86 33.00 11.36 13.91 13.00 13.44 6.82 33.00 11.30 11.58 10.00 10.73 7.22 16.00 9.95
6.2. RQ2: Ablation Study
Table 3 shows the performance of Orienter and its three variations. For Orienter w/o Context Comprehension, we remove the whole Module I Semantic Context Comprehension. For Orienter w/o Reflection-Directed Loop, we remove the whole Module II instead of IGE candidate detection module (to at least localize elements). For Orienter w/o interactability Classification, we remove the whole Module III Context-Sensitive Interactability Classification.
As can be seen in Table 3, Orienter outperforms all its variants significantly, demonstrating the necessity of the designed pipeline. For the Context Comprehension component, its absence results in noticeable performance degradation. Compared to Orienter, the F1 score drops by 14.10% (38.63% vs. 44.97%) and 11.07% (26.62% vs. 29.93%) at 0.75 IoU threshold for the interactability and semantic tasks, respectively. The gap grows even wider as the IoU threshold increases, as such a trend also appears in other metrics, underscoring the importance of semantics context comprehension in improving Orienter’s ability to correctly detect IGEs across varying IoU thresholds. Replacing IGE Feature Mining & Referring components results in a significant weakening of the performance, with an average decline of 19.04% and 12.37% on F1 Score when distinguishing IGEs’ interactability and semantics, suggesting the critical role of the Candidate IGE Referring components. As for the Interactability Classification component, removing this component leads to the most significant performance degradation. Compared to Orienter, the F1 score decreases by 34.60% and 18.94% on average in interactability and semantics, respectively. The experimental results show the necessity of each component in Orienter, highlighting their critical role in accurately detecting IGEs.
6.3. RQ3: Usefulness
(a) Effective interaction count
(b) Coverage rate
Figure 14. Results of RQ3: Usefulness
As shown in Figure 14, the testing guided by Orienter outperforms the random testing with a significant gap. We select the best model, Orienter with Gemini for comparison. The number of effective interactions in guided testing increases notably faster than in random generation. At the end of the testing, the guided one’s average total number of effective interactions reaches 6,571.8, surpassing the random testing’s 2,911.6 by a remarkable 125.7%. Regarding the rate of effective interaction, guided testing peaks at 0.57, with an average of 0.46 over the whole duration of testing, considerably better than that of the random testing remaining at around 0.11. At 10 minutes, the IGE coverage of the guided testing already reaches 0.65, outperforming the random testing’s 0.32 by 103.1%. The IGE coverage of the guided testing eventually reaches 0.83, which is 27.7% higher than the random testing’s 0.65.
These results demonstrate the capability of Orienter to boost the effectiveness and efficiency of VR app testing. Apart from automated VR app testing, other downstream software engineering problems like UI to code, video-based or screenshot-based testing, etc., can also benefit from the ability of Orienter to effective and efficient IGE detection.
7. Threats to Validity
Internal Validity. An internal threat to validity is related to the annotators’ bias and subjective interpretation. Given that the dataset was annotated by 13 annotators, there’s potential bias based on their personal experience and interpretation of interactability in VR apps. This could impact the consistency and accuracy of the annotations. The process of labeling semantics of IGEs is subjective and dependent on the annotator’s understanding and perception, which may not always align with the intended use or perception in different user demographics. To mitigate it, we construct a detailed and objective data collection and annotation process. We spend at least two hours for each annotator to make sure they understand all steps. Another threat is model bias in pretrained models, given that Orienter relies on pretrained LMMs, there is a risk that biases inherent in these models could affect the accuracy of IGE detection. If these models were trained on data that was not representative of diverse contexts, their predictions could be skewed. To mitigate this threat, we choose the most powerful and representative models, which are trained on large corpus and which are claimed that the team has reduced model bias to some level.
External Validity. One threat to external validity is the generalizability across VR apps. The diversity in VR apps poses a challenge to generalizability. Orienter’s effectiveness demonstrated in specific scenarios may not uniformly apply across different VR environments, particularly those with unique interaction paradigms or novel GUI element types. To mitigate this threat, we sample popular and complicated VR apps, crossing a large portion of app categories.
8. Related Work
8.1. Studies on VR/AR Apps
Previous works have conducted several empirical studies to understand VR/AR (XR) apps from different perspectives. Rodriguez and Wang (Rodriguez and Wang, 2017) analyzed the growing trends, amount of developers, popular topics, and common files in open-source VR software projects. Li et al. (Li et al., 2020) conducted an empirical study of bugs in web-based XR projects to understand their symptoms and uniqueness. Adam et al. (Adams et al., 2018) conducted interviews with VR users and developers and surveyed their concerns about the security and privacy of VR apps. Nusrat et al. (Nusrat et al., 2021) systematically performed an empirical study of performance optimization in various VR projects. Li et al. (Li et al., 2023) modelled the software quality attributes and key influencing factors of VR applications from the users’ perspectives. Guo et al. (Guo et al., 2024) developed a security and privacy assessment tool and conducted an empirical study on Oculus VR apps.
8.2. VR/AR Testing
VR/AR (XR) apps provide an immersive experience to users, involving various procedures like device tracking and rendering. To ensure users’ experience, there are several works conducting empirical studies on XR software testing (Andrade et al., 2019). Andrade et al. (Andrade et al., 2019) compared open-source VR projects with non-VR projects to point out the necessity of performing testing on VR apps. Rzig et al. (Rzig et al., 2023) performed a large-scale empirical study on software testing practices of open-source VR projects to identify the current state of testing in VR apps. Several works have proposed approaches to facilitate XR testing. Qin and Hassan (Qin and Hassan, 2022) proposed DyTRec, which provides developers with recommendations on which codes should be tested by actually running the app. Souza et al. (Souza et al., 2018) proposed VR-ReST, a tool designed to assist with requirements specification and test data generation for VR apps. Rafi et al. (Rafi et al., 2022) proposed PredART that can be used as the test oracle when checking the placement of virtual objects in AR. Wang (Wang, 2022) proposed VRTest that extracts information from the VR scene and automatically explores the scene and interacts with the objects by controlling the camera. Andrade et al. (de Andrade et al., 2023) proposed a novel approach that combines metamorphic testing, agent-based testing, and machine learning to test VR apps. Recently, Li et al. (Li et al., 2024) proposed StereoID to automatically detect stereoscopic visual inconsistencies in VR apps.
8.3. User Interface Analysis
Empirical studies have been conducted to give insights into UI analysis (Ye et al., 2021), such as Chen et al. (Chen et al., 2020) conducted a systematic large-scale empirical study on GUI element detection methods. Previous works have adapted both old-fashioned methods  (Yeh et al., 2009; Bao et al., 2015; Dixon and Fogarty, 2010) and deep learning models (Wang et al., 2023a; Qian et al., 2022; Wu et al., 2023) to UI analysis. White et al. (White et al., 2019) replaced methods that involve GUI APIs with machine learning techniques to find interactable elements in GUI images. Xie et al. (Xie et al., 2020) proposed UIED, which combines old-fashioned computer vision approaches and deep learning models to detect components on complex GUI images. For game apps, Ye et al.  (Ye et al., 2021) conducted an empirical study of clickable GUI element detection on mobile games and constructed the first clickable GUI element detection benchmark. Wu et al. (Wu et al., 2023) constructed a larger and more precise GUI dataset for mobile games and evaluated seven GUI detection techniques on it. The industry also applies UI analysis techniques to solve business problems like GUI testing. Ran et al.  (Ran et al., 2022) reported their experiences and lessons learned developing and deploying VTest, an automated visual testing framework for smartphones that leverages only GUI images. Wang et al.  (Wang et al., 2023a) proposed iExplorerGame, a unified framework for black-box mini-game testing that combines deep learning object detection and edge aggregation-based segmentation to guide the testing. Qian et al.  (Qian et al., 2022) proposed a fast OCR-based widget localization technique Label Text Screening, which accelerates the OCR widget detection by analyzing and leveraging the feature of texts in widgets. Research on UI analysis is going in-depth and reaching fields with complex GUI-like game apps. However, existing methods are limited to common IGE with finite predefined categories, which fail to tackle the open vocabulary challenge, as mentioned in  section 1.
9. Conclusion
In this paper, we analyze the key challenges in interactable GUI element (IGE) detection for Virtual Reality (VR) apps. We propose Orienter, the first zero-shot, context-sensitive IGE detection framework for VR apps. Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework. The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, Experiments also illustrate that Orienter is beneficial for boosting automated GUI testing. Alongside developing Orienter, we have also created the first dedicated dataset for IGE detection in VR environments, both of which are publicly available to spur further research in this field.

### v4

@@
-Grounded GUI Understanding for Vision Based Spatial Intelligent Agent: Exemplified by Virtual Reality Apps
+Grounded GUI Understanding for Vision-Based Spatial Intelligent Agent: Exemplified by Extended Reality Apps
 Abstract.
-In recent years, spatial computing Virtual Reality (VR) has emerged as a transformative technology, offering users immersive and interactive experiences across diversified virtual environments.
-Users can interact with VR apps through interactable GUI elements (IGEs) on the stereoscopic three-dimensional (3D) graphical user interface (GUI).
-The accurate recognition of these IGEs is instrumental, serving as the foundation of many software engineering tasks, including automated testing and effective GUI search.
-The most recent IGE detection approaches for 2D mobile apps typically train a supervised object detection model based on a large-scale manually-labeled GUI dataset, usually with a pre-defined set of clickable GUI element categories like buttons and spinners.
-Such approaches can hardly be applied to IGE detection in VR apps, due to a multitude of challenges including complexities posed by open-vocabulary and heterogeneous IGE categories, intricacies of context-sensitive interactability, and the necessities of precise spatial perception and visual-semantic alignment for accurate IGE detection results.
-Thus, it is necessary to embark on the IGE research tailored to VR apps.
-In this paper, we propose the first zero-shot cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps, named Orienter.
-By imitating human behaviors, Orienter observes and understands the semantic contexts of VR app scenes first, before performing the detection.
-The detection process is iterated within a feedback-directed validation and reflection loop.
-Specifically, Orienter contains three components, including (1) Semantic context comprehension for capturing the apps’ GUI context, (2) Reflection-directed IGE candidate detection for identifying and localizing valid GUI elements based on multi-perspective description guided IGE detection, as well as feedback-directed reflection, and (3) Context-sensitive interactability classification which integrates semantic contexts for interactability prediction.
-To evaluate our approach and facilitate follow-up research, we spend more than three months constructing the first benchmark dataset which contains 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories.
-Extensive experiments on the dataset demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches (i.e., GPT-4o, YOLO v8, CenterNet2, Faster R-CNN, UIED, and Xianyu), surpassing their F1 Score by up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) in distinguishing the interactibility and semantics of the IGEs, respectively.
-Orienter is beneficial for boosting the performance of automatic testing by isolating the interactable action space from the whole space, regardless of the testing strategies employed.
-Experiments demonstrate that Orienter-guided testing covers 103.1% more IGEs with 125.7% more effective interactions than testing without action space isolation.
-Virtual Reality, Spatial Intelligence, GUI Agent, Large Multimodal Model
-ccs: Software and its engineering Software development techniquesccs: Computing methodologies Computer visionccs: Human-centered computing Virtual reality
+In recent years, Extended Reality (XR) has emerged as a transformative technology, offering users immersive and interactive experiences across diversified virtual or virtual-real environments.
+Users can interact with XR applications (apps) through interactable GUI elements (IGEs) on the stereoscopic three-dimensional (3D) graphical user interface (GUI).
+IGE constitutes the fundamental element of XR GUI, embodying rich semantic information.
+The accurate recognition and precise understanding of these IGEs is instrumental, serving as the foundation of GUI grounding, which can facilitates downstream tasks, including automated XR testing.
+A straightforward XR test generator can interact randomly within the app’s 3D environment, making it trapped in uninteractable space and resulting in an ineffective and inefficient testing process.
+In contrast, a more intelligent test generator, informed by the accurate locations and semantics of IGEs, can make wiser decisions on interaction targets and orders, forming test sequences that cover more functionalities faster.
+The most recent IGE detection approaches in SE are designed for 2D mobile apps and typically train a supervised object detection model based on a large-scale manually-labeled GUI dataset, usually with a pre-defined set of clickable GUI element categories like buttons and spinners.
+Such approaches can hardly be applied to IGE detection in XR apps, due to a multitude of challenges including complexities posed by open-vocabulary and heterogeneous IGE categories, intricacies of context-sensitive interactability, and the necessities of precise spatial perception and visual-semantic alignment for accurate IGE detection results.
+Thus, it is necessary to embark on the IGE research tailored to XR apps.
+[…]
 1. Introduction
-In recent years, Virtual Reality (VR) has emerged as a transformative technology, offering users immersive experiences across various virtual or virtual-real environments.
-This technological advancement has catalyzed the development of a myriad of VR apps, now numbering in tens of thousands (Li et al., 2023).
+In recent years, Extended Reality (XR) has emerged as a transformative technology, offering users immersive experiences across various virtual or virtual-real environments.
+This technological advancement has catalyzed the development of a myriad of XR apps  (Li et al., 2023).
 These apps, which span a diverse array of domains, including skill training (web, 2021b), entertainment (web, 2021c, a), medical procedures (Khor et al., 2016), and military training (Bhagat et al., 2016), have attracted over 171 million users (Statista, 2022).
 This exponential growth underscores a critical need for robust software processes, including development, testing, and maintenance, particularly in high-reliability contexts such as healthcare and military training.
-In VR apps, users experience multimodal perceptions through various devices and interact via body movements and gestures.
+In XR apps, users experience multimodal perceptions through various devices and interact via body movements and gestures.
 Among the multimodal perceptions, visual perception obtained via an app’s Graphical User Interface (GUI) offers the wealthiest information.
-GUIs in VR apps are often composed of three-dimensional (3D) GUI elements (e.g., images, text, widgets, etc.) or real-life objects and users interact with VR apps through Interactable GUI Elements (short as IGEs).
-Literature has shown that the accurate recognition of IGEs is a cornerstone for many software engineering tasks including automated testing (Xie et al., 2020; Ye et al., 2021; Chen et al., 2020) and effective GUI search (Reiss et al., 2018; Deka et al., 2017).
+GUIs in XR apps are often composed of three-dimensional (3D) GUI elements (e.g., images, text, widgets, etc.) or real-life objects and users interact with XR apps through Interactable GUI Elements (short as IGEs).
+As an important part of GUI grounding to facilitate downstream tasks, literature has shown that the accurate recognition of IGEs is a cornerstone for many software engineering tasks including automated testing (Xie et al., 2020; Ye et al., 2021; Chen et al., 2020) and GUI code generation (Nguyen and Csallner, 2015; Moran et al., 2020; Chen et al., 2018).
 For example, Ye et al.
@@
 (White et al., 2019) demonstrate that IGE detection improves branch coverage by 42.5% compared to random testing.
-Deep learning (DL) based object detection approaches like Faster R-CNN (Ren et al., 2017), CenterNet (Zhou et al., 2021) and YOLO (Jocher et al., 2023), have demonstrated promising performance.
+Deep learning (DL) based object detection approaches (Ren et al., 2017; Zhou et al., 2021; Jocher et al., 2023) have demonstrated promising performance.
 With such advancements, recent work has taken steps to explore DL-based IGE detection approaches in mobile apps and desktop apps (White et al., 2019; Chen et al., 2020; Xie et al., 2020; Ye et al., 2021).
@@
 (Wu et al., 2023) explore advanced DL-based approaches, like CenterNet2, YOLOv3, and YOLOv5, for IGE detection on mobile apps.
-The approaches can be typically divided into three steps: (1) manually summarize a finite set of IGE categories such as buttons, spinners, switches, checkboxes, etc., (2) manually label a large dataset, and then (3) train an object detection model on the dataset.
-However, such training-based approaches heavily rely on large annotated datasets so that they can hardly be directly applied to IGE detection in VR apps.
-As shown in Figure 1(a) and Figure 1(b), annotating large-scale VR IGE datasets is challenging due to the unique interaction mechanism of VR apps.
-The data annotation process demands extensive efforts, compounded by VR’s complex hardware usage and exhaustive interactions.
+The approaches can be typically divided into three steps: (1) manually summarize a finite set of IGE categories such as buttons, input boxes, etc., (2) manually label a large dataset, and then (3) train an object detection model on the dataset.
+However, such training-based approaches heavily rely on large annotated datasets so that they can hardly be directly applied to IGE detection in XR apps.
+As shown in Figure 1(a) and Figure 1(b), annotating large-scale XR IGE datasets is challenging due to the unique interaction mechanism of XR apps.
+The data annotation process demands extensive efforts, compounded by XR’s complex hardware usage and exhaustive interactions.
 Recent advancements of pretrained large multimodal models (LMMs) (OpenAI, 2023; Gemini Team, 2023) have shown their remarkable abilities in downstream tasks including image/natural language comprehension, question answering and logical reasoning, without large training sets on specific problems.
@@
 However, our preliminary experiments (detailed in  section 2.1) reveal that LMMs suffer from severe spatial semantic hallucinations, tending to generate contextually coherent but factually incorrect or unrealistic IGEs with inaccurate locations and amounts as responses.
-These are mainly stemmed from the following challenges of VR IGE detection problem:
+These are mainly stemmed from the following challenges of XR IGE detection problem:
 Challenge #1: Open-vocabulary and heterogeneous GUI element categories.
 LMM or DL approaches perform well on detecting a finite pre-defined set of IGE categories that frequently appear in their training sets and adhere to standard visual appearance patterns (e.g., buttons, sliders, and checkboxes) as well as unified interaction mechanisms (e.g., tapping and long-tapping).
-However, VR apps with diverse scenarios contain infinite open-vocabulary categories of IGEs that own different visual appearances and interaction mechanisms (heterogeneity), as shown in Figure 1(a).
-In the stereoscopic 3D VR scene, from time to time, IGEs occlude each other or are observed from non-front perspectives (e.g., Figure 3(a)), making the visual appearances of even the same IGE overly diverse and heterogeneous.
-Hence, it is hard for LMM or DL approaches to identify such open-vocabulary and heterogeneous VR GUI element categories.
+However, XR apps with diverse scenarios contain infinite open-vocabulary categories of IGEs that own different visual appearances and interaction mechanisms (heterogeneity), as shown in Figure 1(a).
+In the stereoscopic 3D XR scene, IGEs sometimes occlude each other or are observed from non-front perspectives (e.g., Figure 3(a)), making the visual appearances of even the same IGE overly diverse and heterogeneous.
+Hence, it is hard for LMM or DL approaches to identify such open-vocabulary and heterogeneous XR GUI element categories.
 Challenge #2: Context-sensitive interactability.
-The interactability of GUI elements in VR apps is highly dependent on the semantic context of the VR app content and scenarios.
+The interactability of GUI elements in XR apps is highly dependent on the semantic context of the XR app content and scenarios.
 As shown in Figure 1(c) and 1(d), respectively, an object like a tree might be interactable in one app (e.g., a tree-planting game where it needs to be picked up and planted) but non-interactable in another (e.g., as a mere background element in a fishing game).
 The IGE categories that LMM or DL approaches perform well when GUI elements with similar appearances behave mostly the same in distinct contexts, like buttons and checkboxes.
-We call them context-independent GUI elements.
-However, the interactability of GUI elements in VR apps is highly dependent on the app-specific and scenario-specific context, which means they cannot be uniformly determined by appearances but require context-sensitive reasoning instead.
+However, the interactability of GUI elements in XR apps is highly dependent on the app-specific and scenario-specific context, which means they cannot be uniformly determined by appearances but require context-sensitive reasoning instead.
 Challenge #3: Accurate IGE detection results require precise spatial perception and visual-semantic alignment.
@@
 Their ability to outline and label objects is limited, often resulting in imprecise localization, misaligned bounding boxes, etc.
-This spatial deficiency hinders their applicability in VR environments, where precise IGE localization is necessary.
-(a) IGE categories in a VR lab training app (web, 2024d) with diverse visual appearances and interaction mechanisms
-(b) Diversified IGE categories in Figure 1(a) with annotations, dataset annotation requires massive effort
-(c) An interactable tree in a VR gardening (tree planting) game (web, 2024e)
-(d) Several non-interactable trees in a VR fishing game (web, 2024g)
-Figure 1. Examples to demonstrate the challenges of IGE detection for VR apps (web, 2024e, g, d)
-To tackle the challenges, we propose a cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps (short as Orienter).
-Instead of directly identifying the IGEs in VR apps, Orienter is designed to imitate human behavior by comprehending the semantic context first before conducting the detection; and the detection process is iterated within a feedback-directed validation and reflection loop (i.e., two levels of “looking before leaping” in the title).
+This spatial deficiency hinders their applicability in XR environments, where precise IGE localization is necessary.
+(a) IGE categories in lab training app (web, 2024e) with diverse visual appearances and interaction mechanisms
+(b) IGE in diversified categories in Figure 1(a) with annotations, dataset annotation requires massive effort
+(c) An interactable tree in a XR gardening (tree planting) game (web, 2024f)
+(d) Several non-interactable trees in a XR fishing game (web, 2024h)
+Figure 1. Examples to demonstrate the challenges of IGE detection for XR apps (web, 2024f, h, e)
+To tackle the challenges, we propose a cOntext-sensitive inteRactable GUI ElemeNT detection framework for Extended Reality apps (short as Orienter).
+Instead of directly identifying the IGEs in XR apps, Orienter is designed to imitate human behavior by comprehending the semantic context first before conducting the detection; and the detection process is iterated within a feedback-directed validation and reflection loop (i.e., two levels of “looking before leaping” in the title).
 Specifically, Orienter mainly contains three major modules.
-❶ Semantic context comprehension: For identifying heterogeneous IGE categories in an unsupervised and context-sensitive way, we make Orienter to understand and combine the global context (app content overview, including genres, storylines, interaction mechanisms, etc.) and local context (the current VR scenario within the user’s field of view) of the VR scenes under analysis through in-context reasoning of LMM.
+❶ Semantic context comprehension: For identifying heterogeneous IGE categories in an unsupervised and context-sensitive way, we make Orienter to understand and combine the global context (app content overview, including genres, storylines, interaction mechanisms, etc.) and local context (the current XR scenario within the user’s field of view) of the XR scenes under analysis through in-context reasoning of LMM.
 This phase is the foundation of all the subsequent steps.
@@
 ❸ Context-sensitive interactabilitiy classification: To predict the interactibility of detected IGEs, Orienter incorporates the semantic context from the first module and performs chain-of-thought in-context classification.
-For facilitating the evaluation, we spend more than three months constructing the first benchmark dataset for IGE detection in VR apps.
-The benchmark dataset consists of 1,552 images from 100 industrial-setting apps on Steam (web, 2023c), with 4,470 interactable annotations across 766 semantics categories.
-We evaluate Orienter in terms of predictions on interactability and semantics.
-Orienter outperforms all baselines (GPT-4o (web, 2024b), YOLO v8 (Jocher et al., 2023), CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), UIED (Xie et al., 2020), Xianyu (Chen et al., 2020)) in almost all metrics, including precision, recall, and F1 score.
-Specifically, Orienter achieves improvements of up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) on F1 scores in distinguishing the interactability and semantics of IGEs, respectively, demonstrating Orienter’s strong effectiveness, efficiency, and reliability.
-To illustrate the usefulness of Orienter on downstream software engineering (SE) tasks, we further evaluate how effective it boosts the performance of automated testing.
-Orienter’s IGE detection results can be utilized to isolate the interactable action space from the whole input space, regardless of the testing strategies employed.
-Experiment results reveal that Orienter-guided testing covers 103.1% more IGEs with 125.7% more effective interactions than testing without action space isolation.
+[…]
 In summary, we make the following contributions in this paper:
   • To the best of our knowledge, we are the first to formulate the interactable GUI element (IGE) detection problem for stereoscopic 3D GUI.
-We leverage the power of large multimodal models in understanding and analyzing GUIs of VR software, and propose a zero-shot context-sensitive GUI element detection framework, Orienter.
-Orienter novelly undergoes semantic context comprehension, reflection-directed IGE candidate detection, and context-sensitive interactability classification to tackle the challenges of VR IGE detection.
-  • Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework.
-The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, surpassing their F1 Score by up to 3.7$\times$ and 121.4$\times$ (1.4$\times$ and 46.2$\times$ on average) in distinguishing the interactibility and semantics of the IGEs, respectively.
-Experiments also illustrate that Orienter is beneficial for boosting GUI testing by covering 103.1% more IGEs with 125.7% more effective interactions than random testing.
-  • We construct the first dataset for VR IGE detection, including 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories.
-2. Preliminaries
-2.1. Background of Virtual Reality (VR)
-2.1.1. VR Devices
-VR technologies provide immersive experiences that blend virtual elements with the physical world.
-These experiences are accessible via diverse devices including PCs and standalone systems, creating interactive environments that integrate real and virtual elements.
-Key VR devices include: (a) Head-mounted displays (HMDs), essential for VR experiences, offering immersive visuals and audio.
-(b) Two handheld controllers and gesture recognition systems enable users to interact with the virtual world via physical actions and gestures.
-(c) Tracking systems, comprising sensors and cameras, track user movement and orientation, crucial for a responsive VR experience.
-2.1.2. VR Interaction
-Interaction within VR environments can occur through various mechanisms: (a) Pressing buttons on controllers to manipulate GUI elements.
-(b) Touching or moving devices towards GUI elements, sometimes using extended virtual tools for remote interaction.
-(c) Gazing at GUI elements or employing eye-tracking technologies in HMDs to select or move elements.
-(d) Gestures and Movements to interact with the virtual environment through physical actions.
-These interaction modes highlight the evolving nature of human-computer interaction within VR, offering increasingly diverse and immersive experiences.
+We leverage the power of large multimodal models in understanding and analyzing GUIs of XR software, and propose a zero-shot context-sensitive GUI element detection framework, Orienter.
+Orienter novelly undergoes semantic context comprehension, reflection-directed IGE candidate detection, and context-sensitive interactability classification to tackle the challenges of XR IGE detection.
+  • We conduct extensive experiments to verify the effectiveness of the proposed IGE detection framework.
+The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches.
+Experiments show that Orienter can facilitate GUI testing.
+  • We construct the first dataset for XR IGE detection (web, 2025), including 1,552 images from 100 industrial-setting apps on Steam, with 4,470 interactable annotations across 766 semantics categories.
+2. Background on Extended Reality (XR)
+XR Devices.
+XR technologies offer immersive experiences that seamlessly blend virtual and physical worlds.
+These technologies can be accessed through various device types.
+VR typically requires either a PC, a standalone system, or a mobile phone for computational power; AR usually overlays digital information onto the physical world, often through smartphones or specialized AR glasses; and MR integrates the capabilities of both VR and AR, offering more complex and interactive experiences that merge real and virtual environments.
+Typical XR devices include: (1) Head-mounted displays (HMDs), which encompass screens that display digital environments and often with audio capabilities.
+(2) Handheld controllers and gesture recognition systems, which allows users to navigate and manipulate the virtual environment through physical movements.
+(3) Tracking systems, which includes sensors and cameras, tracks user movements and spatial orientation, enabling a responsive and immersive experience in 3D space.
+Figure 2. Prompt template for motivating examples.
+XR Interaction.
+There are various ways in which users can interact with GUI elements.
+Some examples include: (1) Pressing buttons.
+Users can interact with GUI elements by pressing buttons on their devices, such as trigger buttons or joysticks on controllers.
+(2) Touching.
+Users can touch GUI elements by moving their devices to the location of elements in the virtual world.
+Some apps extend users’ hands in the virtual world with a special object to allow users to interact with elements remotely by making the special object contact with others.
+(3) Gazing.
+Users can interact with GUI elements by directing their gaze to them.
+Some headsets support eye tracking that enables interaction using only eye movements.
+(4) Gestures and Movements.
+Users can interact with GUI elements through gestures or movements, such as raising or waving hands.
+As more methods for human-computer interaction are proposed, the interaction methods are becoming more diverse.
+(a) W.r.t heterogeneity
+(b) W.r.t. spatial perception
+Figure 3. LMM gives deviated IGE detection results on VR scenes (web, 2024g, d) because of aforementioned challenges
 2.1. Motivating Examples
 In this section, we present our motivational study to further demonstrate the technical challenges listed in section 1.
-Two authors randomly sample 30 VR apps from the popular Steam VR app store (web, 2023c), select screenshots from them, and use the prompt template in Figure 2 to instruct the widely-used LMM GPT-4o (web, 2024b) to identify and locate IGEs in the images.
-Figure 2. Propmt template for motivating examples.
+Two authors randomly sample 30 VR apps from the popular Steam VR app store (web, 2023c), select screenshots from them, and use the prompt in Figure 2 to instruct LMM GPT-4o (web, 2024c) to identify and locate IGEs in the images.
 As shown in figure 3(a), the mailbox is wrongly recognized as “Missile-like Object”.
@@
 This misunderstanding of GUI elements reveals the LMM’s inability to face the heterogeneous GUI elements in VR apps.
-(a) W.r.t heterogeneity
-(b) W.r.t. spatial perception
-Figure 3. LMM gives deviated IGE detection results on VR scenes (web, 2024f, c) because of aforementioned challenges
 For another example, figure 3(b), the bounding boxes in the image either cover only part of the object, locate objects mismatched with the label, or even mark the background.
 Although the LMM identifies most objects in the image, it fails to locate them well, revealing its insufficient spatial perception.
-To summarize, the LMM alone cannot effectively detect IGEs in highly complex and varied VR apps.
-To boost the LMM on IGE detection in VR apps, it is necessary to provide the LMM with more context information and integrate the model with different techniques.
-Following this idea, we propose Orienter to imitate human behavior to enhance the LMM on detecting IGEs in VR apps.
+To summarize, the LMM alone cannot effectively detect IGEs in highly complex and varied XR apps.
+To boost the LMM on IGE detection in XR apps, it is necessary to provide the LMM with more context information and integrate the model with different techniques.
+Following this idea, we propose Orienter to imitate human behavior to enhance the LMM on detecting IGEs in XR apps.
 3. The Orienter Approach
@@
 Before jumping into the details of our approach, we first introduce how we formulate the IGE detection problem.
-The basic IGE detection unit is an individual VR app scenario, which lies in the user’s field of view, at a specific time.
-Unlike mobile application GUIs or web app GUIs, which are entirely 2D, HMD-based VR apps typically render two 2D GUIs for both eyes of users, creating an illusion of depth and making users feel stereoscopic 3D (S3D) sense (Penn and Hout, 2018; Elmqaddem, 2019).
-These scenes, presented to each eye, are projections of the 3D virtual world onto two 2D ”picture planes” (user’s eyes), mathematically represented by the function: $P_{3D\rightarrow 2D}:UI_{3D}\rightarrow(UI_{left},UI_{right}),$ where $UI_{3D}$ represents the 3D scene in the VR environment, and $UI_{left}$ and $UI_{right}$ represent the corresponding 2D projections rendered for the left and right eyes, respectively.
+The basic IGE detection unit is an individual XR app scenario, which lies in the user’s field of view, at a specific time.
+Unlike mobile application GUIs or web app GUIs, which are entirely 2D, HMD-based XR apps typically render two 2D GUIs for both eyes of users, creating an illusion of depth and making users feel stereoscopic 3D (S3D) sense (Penn and Hout, 2018; Elmqaddem, 2019).
+These scenes, presented to each eye, are projections of the 3D virtual world onto two 2D ”picture planes” (user’s eyes), mathematically represented by the function: $P_{3D\rightarrow 2D}:UI_{3D}\rightarrow(UI_{left},UI_{right}),$ where $UI_{3D}$ represents the 3D scene in the XR environment, and $UI_{left}$ and $UI_{right}$ represent the corresponding 2D projections rendered for the left and right eyes, respectively.
 Each one-eye scene can be recovered using the other eye’s scene (Li et al., 2024).
 Therefore, the S3D IGEs can be detected by analyzing the 2D projection from a single eye.
-This allows us to simplify the VR IGE detection problem to detecting IGEs in the 2D scene of any individual eye.
+This allows us to simplify the XR IGE detection problem to detecting IGEs in the 2D scene of any individual eye.
 To facilitate downstream SE tasks such as boosting automated GUI testing, each IGE is identified not only by its interactability status but also by its location and associated semantic labels, enabling a more fine-grained understanding of the IGEs.
-Let $GE$ represent the set of GUI elements in the VR scene under analysis, i.e., $UI$.
+Let $GE$ represent the set of GUI elements in the XR scene under analysis, i.e., $UI$.
 Given that the 3D scene $UI_{3D}$ is projected into 2D for each eye, the right-eye scene can be denoted as: $P_{3D\rightarrow 2D}(UI_{right}):UI_{3D}\rightarrow UI_{right}.$ The detection function $D:GE_{2D}\rightarrow\{Non-interactable,Interactable\}$ maps each GUI element to a binary interactability determination.
@@
 We use a bounding box $B$ to specify the spatial location of each IGE, which is the smallest unrotated rectangle containing that IGE and can be described by its upper-left corner coordinates, width, and height.
-Let $S$ denote the set of semantic labels associated with each GUI element, where the labels and their granularity depend on the specific VR scene.
+Let $S$ denote the set of semantic labels associated with each GUI element, where the labels and their granularity depend on the specific XR scene.
 As presented above, the problem can be simplified to detection on the 2D scene of any individual eye.
@@
 3.2. Overview of Orienter
-Figure 4 illustrates our cOntext-sensitive inteRactable GUI ElemeNT dEtection framework for virtual Reality apps, short as Orienter.
+Figure 4 illustrates our cOntext-sensitive inteRactable GUI ElemeNT detection framework for Extended Reality apps, short as Orienter.
 To address the three challenges (section 1), Orienter employs a feedback-driven approach that mirrors human behavior, prioritizing semantic comprehension before initiating detection; Orienter also enables iterative refinement and validation through a feedback loop.
 Such designs embody the principle of “looking before leaping”.
-Orienter consists of three primary components: (1) Semantic Context Comprehension: To enable unsupervised, context-sensitive IGE detection, Orienter first synthesizes both the global context (e.g., app genre, storyline, interaction mechanisms) and the local context (i.e., the user’s current VR field of view).
-Leveraging the LMM, Orienter performs in-context reasoning to establish a comprehensive understanding of the VR scene, forming the basis for subsequent detection.
+Orienter consists of three primary components: (1) Semantic Context Comprehension: To enable unsupervised, context-sensitive IGE detection, Orienter first synthesizes both the global context (e.g., app genre, storyline, interaction mechanisms) and the local context (i.e., the user’s current XR field of view).
+Leveraging the LMM, Orienter performs in-context reasoning to establish a comprehensive understanding of the XR scene, forming the basis for subsequent detection.
 (2) Reflection-Guided IGE Candidate Detection: Orienter then mines multi-perspective characteristics of GUI elements, and uses these characteristic descriptions for detecting and locating potential IGEs.
@@
 3.3. Module I: Semantic Context Comprehension
-The interactability of GUI elements in VR apps is highly dependent on the semantic context of the VR app content and scenarios.
-To tackle this challenge, Orienter understands and analyzes the target VR app’s context beforehand to enable context-sensitive detection.
-The semantic context is analyzed comprehensively, from (1) both global (overall app context) and local (current VR scenario) granularities, and (2) both natural language (texts) to vision (images) modalities.
-3.3.1. Global Context Extraction.
-The global context captures the overall semantic information of the VR app under analysis.
-To find out which global context attributes have influences on IGE interactability, two authors randomly sample 30 real-world VR applications from SteamVR and perform manual inspection following the open coding procedure (Creswell and Poth, 2016).
-At last, we reach a consensus and locate the following key attributes: app names, genres, content themes, VR device supports, ways of gameplay, possible interaction mechanisms, and language information.
-Such context can be extracted from the detailed information page on official VR app stores, but effective extraction can be troublesome due to the free structure natural language format of app details.
-Hence, Orienter leverages LMM to perform the extraction after the app details are crawled automatically.
-Prompt PI.1 in Figure 2 shows the prompt template.
+The interactability of GUI elements in XR apps is highly dependent on the semantic context of the XR app content and scenarios.
+To tackle this challenge, Orienter understands and analyzes the target XR app’s context beforehand to enable context-sensitive detection.
+The semantic context is analyzed comprehensively, from (1) both global (overall app context) and local (current XR scenario) granularities, and (2) both natural language (texts) to vision (images) modalities.
 Figure 5.
@@
 Orienter randomly selects three of them as demonstration examples during runtime.
+3.3.1. Global Context Extraction.
+The global context captures the overall semantic information of the XR app under analysis.
+To find out which global context attributes have influences on IGE interactability, two authors randomly sample 30 real-world VR applications from SteamVR and perform manual inspection following the open coding procedure (Creswell and Poth, 2016).
+At last, we reach a consensus and locate the following key attributes: app names, genres, content themes, VR device supports, ways of gameplay, possible interaction mechanisms, and language information.
+Such context can be extracted from the detailed information page on official XR app stores, but effective extraction can be troublesome due to the free structure natural language format of app details.
+Hence, Orienter leverages LMM to perform the extraction after the app details are crawled automatically.
+Prompt PI.1 in Figure 2 shows the prompt template.
 3.3.2. Local Context Perception.
-Local context represents the semantic context within the current VR scenario, i.e., all VR content within the user’s current field of view.
-Under the extracted global context of app content from section 3.3.1, Orienter further guides LMM to perform in-context visual question answering, to digest and summarize the rendered VR scene as local context.
-The guidance of global context ensures that the interpretation of the VR scene is informed by the narrative and purpose of the VR app under analysis.
+Local context represents the semantic context within the current XR scenario, i.e., all XR content within the user’s current field of view.
+Under the extracted global context of app content from section 3.3.1, Orienter further guides LMM to perform in-context visual question answering, to digest and summarize the rendered XR scene as local context.
+The guidance of global context ensures that the interpretation of the XR scene is informed by the narrative and purpose of the XR app under analysis.
 Specifically, the LMM processes the composite input (screenshot and global context) and outputs a summary of all GUI elements and the background within the scene.
@@
 3.4. Module II: Reflection-Directed IGE Candidate Detection
-This module aims at recognizing and localizing IGE candidates within the VR scene, based on the captured semantic contexts from Module I.
+This module aims at recognizing and localizing IGE candidates within the XR scene, based on the captured semantic contexts from Module I.
 LMM can reveal more semantic hints than traditional methods, making it a better backbone for our approach.
 However, LMM tends to generate semantically-incorrect IGEs with inaccurate locations.
-This is because VR IGE detection requires precise spatial perception and visual-semantic alignment capabilities, and VR IGE categories are open-vocabulary and heterogeneous (as presented in the challenges and motivational examples).
+This is because XR IGE detection requires precise spatial perception and visual-semantic alignment capabilities, and XR IGE categories are open-vocabulary and heterogeneous (as presented in the challenges and motivational examples).
 Orienter conquer these challenges through a reflection-directed IGE candidate detection loop:
@@
 Adding more descriptions about the target element (like color, size, shape, relevant locations, etc.) makes the detection more effective and robust.
-It is similar for To boost the effectiveness and robustness of i) and ii), Orienter extends a direct LMM detector to a multi-perspective characteristics miner (section 3.4.1).
+To boost the effectiveness and robustness of i) and ii), Orienter extends a direct LMM detector to a multi-perspective characteristics miner (section 3.4.1).
 Overall, Orienter first mines multi-perspective characteristics of GUI elements leveraging LMM, and then uses these characteristic descriptions as guidance to detect corresponding elements.
@@
 3.4.1. Multi-Perspective Characteristics Mining
-Figure 6. Example input for Module II.①, from the Baseball Kings VR app (web, 2024a)
-In this module, Orienter firstly identifies all IGE candidates in the VR scene, and then mines their diverse characteristics from different perspectives (e.g., size, color, shape, relevant location with other objects, etc.).
+In this module, Orienter firstly identifies all IGE candidates in the XR scene, and then mines their diverse characteristics from different perspectives (e.g., size, color, shape, relevant location with other objects, etc.).
 Then these characteristics will form a characteristics description, which can uniquely identify the corresponding IGE candidate in the present figure in the described detection module.
@@
 Figure 6. Prompt templates and examples for Module II.①2
+Figure 7. Example input for Module II.①, from the Baseball Kings VR app (web, 2024a)
 II.①(a): IGE Candidate Recognition (Prompt PII.1).
-Orienter engages in the recognition of IGE candidates within the VR scene.
+Orienter engages in the recognition of IGE candidates within the XR scene.
 With the semantic context from Module I (section 3.3), Orienter analyzes the visual semantic cues presented in the environment, finding out all IGE candidates.
@@
 Following the identification of IGE candidates, Orienter embarks on recognizing the characteristics dimensions of them, which makes them distinguishable from other GUI elements.
-This involves dissecting the unique characteristics attributes and properties related to the visual appearance (e.g., color, size, shape, etc.), functionalities, interaction mechanisms, and relative locations of each GUI element.
+This involves dissecting the unique characteristics attributes and properties related to the visual appearance, functionalities, interaction mechanisms, and relative locations of each element.
 II.①(c) Characteristics Question Formulation (Prompt PII.3).
@@
 II.①(d) In-Context Characteristics Reasoning (Prompt PII.4).
-Then Orienter conducts characteristics reasoning, querying the LMM to analyze the VR scene and find out the answers for formulated characteristic questions.
-This step integrates the local and global contexts previously established, allowing the framework to interpret the responses within the broader narrative and functional scope of the VR application.
+Then Orienter conducts characteristics reasoning, querying the LMM to analyze the XR scene and find out the answers for formulated characteristic questions.
+This step integrates the local and global contexts previously established, allowing the framework to interpret the responses within the broader narrative and functional scope of the XR application.
 Through this in-context reasoning, Orienter synthesizes a comprehensive understanding of each IGE candidate’s characteristics.
@@
 3.4.2. Described IGE Candidate Detection
-Figure 8. The framework of described IGE candidate detection module (VFMD)
-Figure 9. Example: described IGE candidate detection under the CD “donut with colored granules on the surface”, the donut in the middle is detected.
+Figure 8. The framework of the described IGE candidate detection module (VFMD)
 Orienter detects and localizes IGEs using the descriptions of the crafted characteristics generated in the previous step, inspired by Shen’s work (Shen et al., 2024).
@@
 Figure 8 shows the framework structrure.
-VFMD consists of a vision backbone for extracting visual features from VR scenes and a language model for generating text embeddings from CDs.
+Figure 9. Example: described IGE candidate detection under the CD “donut with colored granules on the surface”, the donut in the middle is detected.
+VFMD consists of a vision backbone for extracting visual features from XR scenes and a language model for generating text embeddings from CDs.
 Text embedding s are aggregated into sentence-level representations and are fused with visual embeddings using a cross-modality transformer encoder.
@@
 A final visual-language alignment module predicts the correct pairings of IGE candidate locations and CDs.
-The backbone language model, trained on a vast corpus of data, can generalize to the diverse and unlimited categories of GUI elements present in VR scenarios.
-The localization process cross-references (or aligns with) the textual CDs with the visual input from the VR scene.
-By doing so, VFMD successfully locates IGE candidates.
-Here is an example to demonstrate the output of this module, as shown in Figure 9.
+The backbone language model, trained on a vast corpus of data, can generalize to the diverse and unlimited categories of GUI elements present in XR scenarios.
+The localization process cross-references (or aligns with) the textual CDs with the visual input from the XR scene.
+By doing, VFMD successfully locates IGE candidates.
+Figure 9 is an example showing the output of this module.
 3.4.3. Feedback-Directed Reflection
@@
 Conversely, for successfully detected IGEs, we design a method of regional zooming and mirroring.
-This involves cropping the detected GUI element within its bounding box into a new figure, and Orienter lets LMM conduct a comparative analysis against the original VR scene.
+This involves cropping the detected GUI element within its bounding box into a new figure, and Orienter lets LMM conduct a comparative analysis against the original XR scene.
 Such an approach facilitates a granular verification of the detection’s accuracy.
@@
 Figure 11. Prompt templates for Module III2
-After establishing the presence and location of GUI elements in the VR environment through the first two modules, Orienter proceeds to analyze their interactivity.
-This module determines which GUI elements in the VR scene can be interacted with.
+After establishing the presence and location of GUI elements in the XR environment through the first two modules, Orienter proceeds to analyze their interactivity.
+This module determines which GUI elements in the XR scene can be interacted with.
 Such information is useful for many downstream SE tasks, including boosting automated testing performance (detailed in section 6.3).
@@
 To classify interactability, Orienter guides LMM towards a chain-of-thought in-context reasoning process, within the semantic context captured in Module I.
-This process mimics a VR player’s thought pattern in distinguishing between GUI elements that are mere visual elements and those that afford interaction.
+This process mimics a XR player’s thought pattern in distinguishing between GUI elements that are mere visual elements and those that afford interaction.
 For instance, a tree in a tree-planting game (Figure 1(c)) is interactable, serving as a GUI element to be manipulated by the player, whereas the same tree would be non-interactable background scenery in a fishing game (Figure 1(d)).
@@
 4. Dataset Construction
-To the best of our knowledge, there is no existing GUI dataset for IGE detection in VR apps. Therefore, we build a dataset to verify the usefulness and effectiveness of Orienter.
-We recruit a team of 13 annotators, spending more than three months in the collection and annotation of GUI images from various VR apps.
+To the best of our knowledge, there is no existing GUI dataset for IGE detection in XR apps. Therefore, we build a dataset to verify the usefulness and effectiveness of Orienter.
+We recruit a team of 13 annotators, spending more than three months in the collection and annotation of GUI images from various XR apps.
 These annotators have a minimum of two years of computer science or electrical engineering background, and mostly with experience in video games.
-The annotators first interact with GUI elements in various VR apps while their views are recorded.
+The annotators first interact with GUI elements in various XR apps while their views are recorded.
 They then select GUI images from these recordings and annotate IGEs.
-To enhance the quality of the dataset, we train the annotators in the use of VR equipment and the labeling of IGEs, and we ask them to follow several guidelines during data collection and annotation: (1) attempt to identify all IGEs using every possible interaction method in each VR scene, and (2) categorize the semantics of IGEs with appropriate granularity based on their context.
-4.1. Collection of VR Apps
-We first collect VR apps from the Steam app store (web, 2023c), a comprehensive repository with a wide variety of VR content.
+To enhance the quality of the dataset, we train the annotators in the use of XR equipment and the labeling of IGEs, and we ask them to follow several guidelines during data collection and annotation: (1) attempt to identify all IGEs using every possible interaction method in each XR scene, and (2) categorize the semantics of IGEs with appropriate granularity based on their context.
+4.1. Collection of XR Apps
+We first collect XR apps from the Steam app store (web, 2023c), a comprehensive repository with a wide variety of XR content.
 This collection yielded a total of 4,610 VR-Only apps, which includes software (official type label on Steam referring to non-games) and games.
 To obtain a statistically significant sample size with a 95% confidence level and a 10% margin of error, we randomly sample 102 apps for analysis to cover a broader spectrum of categories, contexts, and interaction paradigms, covering 245 community-generated genres on Steam, demonstrating their diversity and representativeness.
-4.2. Collection of GUI Images from VR Apps
-In this step, the annotators engage with the selected VR apps, interacting with all GUI components within the VR scene in every possible way, as described in section 2.1.2, while recording their stereo view.
-The annotators then select and save GUI images that encompass all VR scenes and GUI elements they explored from the recordings.
-We crop the right-eye images from the stereo-view recordings, as only one side of the view is needed, as explained in section 3.1.
+4.2. Collection of GUI Images from XR Apps
+In this step, the annotators engage with the selected XR apps, interacting with all GUI components within the XR scene in every possible way, as described in fig.
+2, while recording their stereo view.
+The annotators then select and save GUI images that encompass all XR scenes and GUI elements they explored from the recordings.
+We crop the right-eye images from the stereo-view videos as only one side of the view is needed, as explained in section 3.1.
+4.3. Annotation of IGEs in GUI Images
 (a) Cooking (web, 2023d)
@@
 Figure 12. Examples to demonstrate varying granularity to best categorize semantics of IGEs.
-4.3. Annotation of IGEs in GUI Images
-In this step, the annotators identify IGEs in GUI images and label their locations with bounding boxes as well as their semantics within the context.
-As discussed in section 1, the interactability and semantics of GUI elements heavily depend on the diverse contexts of VR apps, making it challenging to categorize all GUI elements with a finite set of predefined categories.
+The annotators then identify IGEs in GUI images and label their locations with bounding boxes as well as their semantics within the context.
+As discussed in section 1, the interactability and semantics of GUI elements heavily depend on the diverse contexts of XR apps, making it challenging to categorize all GUI elements with a finite set of predefined categories.
 To address this challenge, we ask the annotators to create categories.
-Specifically, they are instructed to classify IGEs with appropriate granularity based on their experience, to better describe the semantics of IGEs in concrete contexts.
-For example, in the VR game, VR The Diner Duo (web, 2023d), fish is only one kind of ingredient to make burgers, as shown in 12(a).
+Specifically, the annotators are instructed to classify IGEs with appropriate granularity based on their experience to better describe the semantics of IGEs in concrete contexts.
+For example, in the XR game, VR The Diner Duo (web, 2023d), fish is only one kind of ingredient to make burgers, as shown in 12(a).
 Different individuals of fish are semantically the same in this game.
@@
 4.4. Dataset Statistics
-After data cleaning, we finally construct a dataset consisting of 1,552 images with the size of 960 * 540 from 100 apps, covering 245 community-generated genres on Steam, with 4,470 interactable annotations across 766 semantics categories, highlighting the diversity of our dataset.
+After data cleaning, we construct a dataset consisting of 1,552 images with the size of 960 * 540 from 100 apps, covering 12 official app genres and 245 community-generated app categories on Steam, with 4,470 interactable annotations across 766 semantics categories, highlighting the diversity of our dataset.
 Annotations are in widely used COCO (Lin et al., 2014) format, which uses bounding boxes to mark the locations of IGEs.
+Detailed distribution for apps in app genres and annotations in IGE’s semantics categories are shown in Figure 13(a) and Figure 13(b), respectively.
+(a) # App in different genres. Genres in different color are used as different set in genre splitting of the dataset introduced in section 5.3.2
+(b) # Annotation in the First 30 IGE categories associated with the most annotations. The IGE category ”button” which contributes 555 annotations is ommitted
+Figure 13. Dataset distribution
 5. Experiment Design
@@
 In this study, our experiment is designed to answer the following research questions:
-  • RQ1 (Performance in industrial-setting): How effective is our proposed framework, Orienter, in IGE detection on industrial-setting VR apps?
+  • RQ1 (IGE detection performance in industrial-setting): How effective is Orienter in IGE detection on industrial-setting XR apps?
     • RQ1-1 (Interactability): How effective is Orienter in terms of analyzing interactability?
     • RQ1-2 (Semantics): How effective is Orienter in terms of inferring semantics?
-    • RQ1-3 (Context-sensitive interactability): How does Orienter perform in terms of analyzing context-sensitive interactability?
-  • RQ2 (Ablation Study): How does each component of Orienter contribute to its performance?
-  • RQ3 (Usefulness): How effectively can Orienter boost automated testing on VR apps?
+    • RQ1-3 (Context-sensitive interactability): How does Orienter perform in terms of analyzing context-sensitive interactability, i.e.
+the same kind of object with different interactability in different contexts?
+  • RQ2 (Ablation Study): How does each component of Orienter contributes to its performance?
+  • RQ3 (Effectiveness for improving testing): How effectively can Orienter boost automated testing on XR apps?
 5.2. Baselines
 Figure 14. Direct prompting
-GPT-4o (web, 2024b) & Gemini 1.5 Pro (Reid et al., 2024) (direct prompting): Simply prompt the model with the image and ask it to locate all IGEs with bounding boxes and their semantics in the image to illustrate the end-to-end performance of LMM.
-Xianyu (Chen et al., 2020) is a UI-to-Code tool developed by Alibaba, leveraging old-fashioned computer vision techniques and OCR.
-UIED (Xie et al., 2020) combines old-fashioned methods with deep learning models to detect clickable GUI elements in complex GUI images.
-CenterNet2 (Zhou et al., 2021) is a two-stage object detector that uses class-agnostic one-stage detectors as the proposal network.
-It estimates object probabilities in the first stage and conditionally classify objects in the second stage.
-Faster R-CNN (Ren et al., 2017) is a two-stage anchor-based deep learning object detector.
-It contains a region proposal network to extract regions of interest.
-Objects within the RoIs are then classified with another neural network.
-YOLO v8 (Jocher et al., 2023) is a one-stage anchor-free detection model.
-It detects and classifies objects simultaneously and thus is faster than two-stage detectors.
+GPT (web, 2024c) & Gemini (Reid et al., 2024) & Claude (web, 2024b) (direct prompting): Simply prompt the model with the image and ask it to find the location and semantics of all IGEs.
+Xianyu (Chen et al., 2020) is a UI-to-Code tool leveraging old-fashioned computer vision techniques and OCR.
+UIED (Xie et al., 2020) combines old-fashioned methods with deep learning models to detect clickable elements in GUI images.
+CenterNet2 (Zhou et al., 2021) is a two-stage object detector that first estimates object probabilities images and then conditionally classify objects.
+Faster R-CNN (Ren et al., 2017) is a two-stage anchor-based deep learning object detector with a region proposal network and a classification network.
+YOLO v8 (Jocher et al., 2023) is a one-stage anchor-free detection model that detects and classifies objects simultaneously.
+OmniParser (Wan et al., 2024) combines object detection, OCR, and LMM to parse GUI images into element trees for GUI grounding.
 5.3. Implementation Details and Experimental Setup
@@
 For described IGE detection, we leverage the pretrained model of APE-L_D (Shen et al., 2024), which is trained on ten datasets and demonstrates promising results in visual grounding.
-We apply post-processing to improve Orienter’s prediction quality by filtering abnormally large bounding boxes (over 90% of the image size), and applying Non-Maximum Suppression (NMS) which retains only the highest-confidence box among overlapping ones with IoU exceeds 0.7 to reduce duplicates.
+We apply post-processing to improve Orienter’s prediction quality by filtering abnormally large bounding boxes ( $90\%$ image size), and applying Non-Maximum Suppression (NMS) which retains only the highest-confidence box among overlapping ones with IoU exceeds 0.7 to reduce duplicates.
 5.3.2. Dataset Preparation
@@
 (3) Context dataset: Contains 41 categories randomly sampled from the most common 100 categories, and extra annotations marking their corresponding non-interactable objects, testing methods’ understanding of context-sensitive IGEs in different contexts, answering RQ1-3.
-We partitioned the images into training/validation/testing datasets in a 6:1:3 ratio using three distinct methods: App split, Genre split, and Context-sensitive split.
-The App split randomly allocates images based on their corresponding apps.
-The Genre split further considers the apps’ genres to assess the methods’ performance across different app types.
-The Context-sensitive split assigns images containing the 41 sampled categories in the context dataset to the test set while randomly distributing the rest into the training and validation sets.
+Table 1. Statistics of different splitting methods. Ctx. is short for Context.
+Splitting Method Set # Covered App # GUI Scene Image # IGEs
+- Full 100 1,552 4,470
+App Train 60 925 2,459
+Valid 10 163 574
+Test 30 464 1,437
+Genre Train 65 958 2,528
+Valid 11 136 610
+Test 24 458 1,332
+Ctx.-sensitive Train 96 947 2,583
+Valid 64 158 404
+Test 53 447 1,483
+Test
+(Ctx. dataset)
+53 447 1,239
+$\bigcup${App, Genre} Test 46 687 2,034
+$\bigcup${Ctx.-sensitive,
+App, Genre}
+Test 77 936 2,875
+We partition the images into training/validation/test sets in a ratio of approximately 6:1:3 using three distinct splitting methods: App splitting, Genre splitting, and Context-sensitive splitting.
+App splitting randomly allocates images based on their corresponding apps such that images from the same app are placed in the same set.
+Genre splitting further considers the apps’ genres to assess the methods’ performance across different app types.
+Context-sensitive splitting assigns images containing the 41 sampled categories in the context dataset to the test set while randomly distributing the remaining images into the training and validation sets.
+The statistics of each splitting method are illustrated in Table  1. The usage of the datasets and splitting methods will be discussed in section 5.3.3 and section 5.3.4.
 5.3.3. Experimental Setup
-We train baselines on the training dataset and evaluate all methods including Orienter on the test dataset.
+We train baselines on the training set and evaluate all methods including Orienter on the test set.
 For Xianyu (Chen et al., 2020), we only use its component detection part to find IGEs.
-For UIED (Chen et al., 2020), we reimplement and train the CNN classifier on our dataset to adapt it to our problem, and replace the OCR component with PaddleOCR.
-For training on CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), and YOLO v8 (Jocher et al., 2023), we set the batch size to 32 and accordingly adjust the learning rate linearly, and keep other configurations the same as their original releases.
-We apply early stopping with the patience of 20 epochs to avoid overfitting.
-For LMM experiments, we run three times and regard the average results as final results.
-We implement a semantic matching tool to support open vocabulary category matching.
+For UIED (Chen et al., 2020), we train the CNN classifier on our dataset to adapt it to our problem.
+For training on CenterNet2 (Zhou et al., 2021), Faster R-CNN (Ren et al., 2017), and YOLO v8 (Jocher et al., 2023), we set the batch size to 32 and accordingly adjust the learning rate, apply early stopping with the patience of 20 epoch.
+Other configurations are the same as their original releases.
+For LMM experiments, we run three times and regard the average results as the final results.
+In terms of evaluation, we implement a semantic matching tool to support open vocabulary category matching.
 We consider two categories semantically match if the cosine similarity of their embedding vectors obtained from the embedding-3 model released by Zhipu AI (Zhipu AI, 2024) exceeds a preset threshold.
-The first two authors examine the similarity of some typical categories and set the threshold as 0.85.
-For RQ1-1, RQ1-2, and RQ2, predictions are evaluated with our customized COCO API that adapts to our semantic matching tool.
-Note that during evaluation, instead of simply ignoring categories without ground truth annotations (usually due to splitting the dataset), as the official COCO API (web, 2023a) does, we further check if any predictions fall into those categories.
-Metrics of categories that meet this condition are set to 0 and included when averaging metrics across categories, providing more comprehensive results.
-Regarding RQ1-3, the baseline methods are trained on the Semantic dataset using Context-sensitive Split.
-Predictions on the test set are then evaluated on the Context dataset with the same split.
+The first two authors examine the similarity of some typical categories and set the threshold to 0.85.
+We accrodingly modify the official COCO API (web, 2023a) to support semantic matching.
+For RQ1-1, RQ1-2, and RQ2, predictions are evaluated with the modified COCO API.
+We leverage App splitting and Genre splitting to evaluate each method in RQ1-1 and RQ1-2, then use the union of both splitting’s test set to assess the contribution of each components of Orienter in RQ2.
+Regarding RQ1-3, the baseline methods are trained on the Semantic dataset using Context-sensitive Splitting.
+Predictions on the test set are then evaluated on the Context dataset with the same splitting.
 The more IGEs and less non-interactable objects the method detects, the better this method’s performance on context-sensitive interactability understanding.
-Metrics are initially calculated for each category and then averaged to obtain the final results.
-5.3.4. VR App Automated Testing Setup
-To answer RQ3, we simulate a simplified test scenario: given a screenshot of the scene in the VR app under test, the testing agent attempts to interact with the GUI elements in the scene in a black-box setting.
-We compare the performance of the testing agent with and without the guidance of Orienter using the test set of Genre split as input.
-We simplify the interaction events in the 3D virtual space to interaction points on UI images of VR scenes and ignore the type of interactions.
+5.3.4. XR App Automated Testing Setup
+To answer RQ3, we simulate a simplified test scenario: given a screenshot of the scene in the XR app under test (AUT), the testing agent attempts to interact with the GUI elements in the scene in a black-box setting.
+We compare the performance of the testing agent with and without the guidance of Orienter using the union of test sets of the 3 splitting as input, regarding the covered 77 apps as AUTs.
 The testing process is modeled as incrementally generating points on UI images over time, treating each point as an interaction made by the testing agent with an interval of one minute.
-We follow the recent works (Pan et al., 2020; Ghorbani et al., 2023) to set the testing duration to 60 minutes, which is longer than exploring a VR scene usually requires.
-We perform 5 testing runs and average the metrics to produce comprehensive results for each interaction strategy.
-For the non-guided strategy, the testing agent randomly generates interaction points on the whole image.
-For the guided strategy, in each attempt, the testing agent generates interaction within the bounding boxes predicted by Orienter at a probability $p$, or ignores the prediction’s constraint to randomly explore the whole UI images at the probability $(1-p)$.
+We follow the recent works (Pan et al., 2020; Ghorbani et al., 2023) to set the testing duration to 60 minutes, which is longer than exploring a XR scene usually requires.
+We perform 30 testing runs and average the metrics for each interaction strategy.
+For the non-guided (i.e.
+random) strategy, the testing agent randomly generates interaction points on the whole image.
+For the guided strategy, in each attempt, the testing agent generates interaction within the bounding boxes predicted by Orienter at a probability $p$, or randomly explore the whole UI images at the probability (1p)(1-p).
 The probability $p$ decreases gradually as the test progresses.
-Suppose the current moment is $t$ minutes and the total duration of the test is $T$ minutes, then $p=1-t/T$.
+Suppose the current moment is $t$ minutes and the total duration of the test is $T$ minutes, then p=1t/Tp=1-t/T.
 This strategy allows the agent to generate effective yet diversified interaction events to cover more IGEs precisely.
@@
 Precision, Recall, and F1-Score.
-A predicted bounding box matches the one with the highest IoU among the ground-truth bounding boxes that (1) lie in the same image, (2) have matching categories, and (3) have IoU exceeding the threshold.
+A predicted bounding box matches the one with the highest IoU among the ground-truth bounding boxes that (1) have matching categories, and (2) have IoU exceeding the threshold.
 For RQ1-1, RQ1-2, and RQ2, we consider a predicted box matching a ground-truth box a True Positive (TP); otherwise, a False Positive (FP); ground-truth boxes that fail to match any predicted box are False Negatives (FN).
-For RQ1-2, a predicted box is considered a TP if it matches an interactable ground-truth box, and an FP if it matches a non-interactable one; non-interactable and interactable ground-truth boxes that fail to match any predicted box are considered a TN and FN, respectively.
-We calculate Precision as $TP/(TP+FP)$, Recall as $TP/(TP+FN)$ and F1-Score as $2\times Precision\times Recall/(Precision+Recall)$.
+For RQ1-2, a predicted box is considered a TP if it matches an interactable ground-truth box, and an FP if it matches a non-interactable one; non-interactable and interactable ground-truth boxes that fail to match any predicted box are considered as TNs and FNs, respectively.
+We calculate $Precision=TP/(TP+FP)$, $Recall=TP/(TP+FN)$ and $F1=2\times Precision\times Recall/(Precision+Recall)$.
 AP and mAP.
@@
 The mean AP (mAP) is obtained by averaging AP across categories.
-We follow the previous works (Wang, 2022; Wang et al., 2023b) to use IGE Coverage and Effective Interaction Count in RQ3.
-Note that we can not calculate code coverage metrics in the black-box setting of RQ3.
+We follow previous works (Wang, 2022; Wang et al., 2023b) to use IGE Coverage and Effective Interaction Count in RQ3.
 IGE Coverage.
-The IGE coverage reflects the effectiveness of interaction strategies to find IGEs over time.
-An IGE is covered if at least one interaction point falls within the bounding box of that IGE.
+The IGE coverage reflects the effectiveness of strategies to find IGEs over time.
+An IGE is covered if at least one interaction point falls in the bounding box of that IGE.
 IGE Coverage is calculated as $n_{covered}/n_{all}$, where $n$ represents the number of IGEs.
 Effective Interaction Count.
-The Effective Interaction Count reflects the efficiency of interaction strategies to make effective interactions by focusing on IGEs only.
+The Effective Interaction Count reflects the efficiency of strategies to make effective interactions by focusing on IGEs only.
 An interaction point is considered effective if it falls in any of IGEs’ bounding boxes.
 6. Results and Analysis
-6.1. RQ1: Performance in industrial-setting
+6.1. RQ1: IGE detection performance in industrial-setting
+Table 2. Performance of Orienter, DL-based baselines are trained on our dataset*
+Methods Interactability Semantics
+App Splitting Genre Splitting App Splitting Genre Splitting
+IoU 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95
+Faster-
+RCNN
+mAP% 18.05 13.84 7.32 1.47 0.07 17.90 12.88 6.96 1.56 0.13 0.21 0.15 0.03 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+P% 33.41 32.29 23.04 11.08 2.15 33.42 29.55 20.10 8.06 1.14 0.91 0.83 0.38 0.16 0.06 0.07 0.06 ¡0.01 ¡0.01 ¡0.01
+R% 32.00 25.00 18.00 8.00 1.00 29.00 22.00 15.00 6.00 1.00 4.00 4.00 3.00 3.00 4.00 2.00 2.00 ¡0.01 ¡0.01 ¡0.01
+F1% 32.69 28.18 20.21 9.29 1.36 31.05 25.22 17.18 6.88 1.06 1.48 1.37 0.68 0.30 0.11 0.13 0.12 ¡0.01 ¡0.01 ¡0.01
+Center-
+Net2
+mAP% 17.79 14.98 11.73 5.91 0.93 15.63 12.53 9.07 4.23 0.45 0.20 0.18 0.15 0.05 ¡0.01 0.22 0.19 0.05 0.02 ¡0.01
+P% 32.53 28.52 26.85 18.26 4.83 28.67 25.81 27.71 16.46 4.51 0.68 0.65 0.57 0.27 0.05 0.62 0.59 0.27 0.19 0.04
+R% 32.00 29.00 23.00 15.00 3.00 30.00 27.00 18.00 12.00 3.00 4.00 4.00 3.00 3.00 2.00 5.00 5.00 3.00 3.00 3.00
+F1% 32.26 28.76 24.77 16.47 3.70 29.32 26.39 21.82 13.88 3.60 1.16 1.12 0.96 0.49 0.10 1.10 1.06 0.49 0.35 0.08
+Yolo v8 mAP% 15.30 13.12 9.85 6.16 1.51 11.54 10.02 7.59 3.86 0.46 0.02 0.02 0.01 ¡0.01 ¡0.01 0.03 0.02 0.02 0.01 0.01
+P% 26.58 25.38 20.63 16.65 7.31 23.92 21.38 24.22 14.55 2.68 0.21 0.20 0.19 0.05 0.01 0.29 0.28 0.27 0.25 0.21
+R% 26.00 23.00 20.00 14.00 3.00 23.00 22.00 15.00 11.00 3.00 4.00 4.00 4.00 1.00 3.00 3.00 3.00 3.00 3.00 3.00
+F1% 26.29 24.13 20.31 15.21 4.25 23.45 21.69 18.53 12.53 2.83 0.40 0.38 0.36 0.09 0.02 0.53 0.52 0.50 0.46 0.40
+OmniParser mAP% 0.13 0.13 0.13 0.05 ¡0.01 0.29 0.28 0.21 0.08 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+P% 2.48 2.44 2.22 2.37 0.29 6.19 5.90 2.93 1.76 0.96 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+R% 2.00 2.00 2.00 1.00 ¡0.01 3.00 3.00 3.00 2.00 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+F1% 2.21 2.20 2.10 1.41 ¡0.01 4.04 3.98 2.96 1.87 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+Gemini 1.5
+Pro
+mAP% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.11 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.12 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+P% 0.21 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.43 0.43 ¡0.01 ¡0.01 ¡0.01 0.25 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.23 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+R% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 4.00 ¡0.01 ¡0.01 ¡0.01 ¡0.01 50.00 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+F1% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.47 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.46 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+gpt-4o-
+2024-08-06
+mAP% 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+P% 0.78 ¡0.01 ¡0.01 ¡0.01 ¡0.01 0.75 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+R% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+F1% ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+Orienter
+(gpt-4-vision-
+preview)
+mAP% 25.53 24.32 23.03 19.64 4.61 27.07 25.22 23.87 19.80 4.54 14.93 14.66 13.96 12.17 3.67 15.96 15.57 14.62 12.87 3.75
+P% 34.79 33.52 32.32 29.31 14.92 38.42 36.70 34.97 31.40 16.79 16.06 16.05 15.17 15.27 6.53 17.11 16.85 15.78 15.93 6.14
+R% 40.00 39.00 38.00 34.00 12.00 40.00 38.00 37.00 32.00 11.00 50.00 50.00 50.00 33.00 20.00 50.00 50.00 50.00 33.00 20.00
+F1% 37.21 36.05 34.93 31.48 13.30 39.20 37.34 35.95 31.70 13.29 24.31 24.30 23.28 20.88 9.84 25.50 25.21 23.98 21.48 9.39
+Orienter
+(gpt-4o-
+2024-08-06)
+mAP% 26.44 25.26 23.98 20.35 4.48 26.11 24.50 23.51 19.42 4.38 18.00 17.65 16.91 14.28 5.75 19.45 18.97 18.07 15.40 5.68
+P% 34.63 34.89 35.43 31.64 18.57 37.91 36.68 36.71 33.24 12.91 19.32 18.97 18.28 15.37 6.87 20.35 19.81 19.01 16.19 7.36
+R% 42.00 39.00 36.00 34.00 11.00 41.00 39.00 37.00 33.00 14.00 50.00 50.00 50.00 50.00 42.00 50.00 50.00 50.00 50.00 25.00
+F1% 37.96 36.83 35.71 32.78 13.81 39.39 37.80 36.85 33.12 13.43 27.87 27.50 26.77 23.52 11.81 28.93 28.38 27.55 24.46 11.37
+Orienter
+(Claude 3.5
+Sonnet)
+mAP% 23.99 22.90 21.54 18.46 3.90 25.50 23.96 22.21 18.49 3.84 16.14 15.79 15.17 12.73 5.01 19.01 18.51 17.78 15.00 6.67
+P% 31.73 37.52 36.79 33.56 16.30 37.97 40.15 39.22 35.86 14.51 16.63 16.28 15.83 13.35 7.64 19.68 19.11 18.53 15.75 9.87
+R% 40.00 32.00 31.00 29.00 11.00 37.00 33.00 32.00 29.00 13.00 50.00 50.00 50.00 50.00 20.00 50.00 50.00 50.00 50.00 20.00
+F1% 35.39 34.54 33.65 31.11 13.14 37.48 36.23 35.24 32.07 13.71 24.95 24.56 24.04 21.07 11.06 28.24 27.66 27.04 23.96 13.22
+Orienter
+(Gemini 1.5
+Pro)
+mAP% 32.71 31.27 30.51 26.33 5.88 35.03 32.59 30.82 25.08 5.64 20.69 20.35 20.19 17.77 7.16 22.12 21.44 21.17 18.45 7.41
+P% 46.42 45.30 44.94 41.64 21.46 49.45 47.52 46.42 41.94 19.02 22.64 22.29 22.16 19.39 9.66 23.72 22.86 22.64 19.58 8.35
+R% 42.00 41.00 40.00 37.00 14.00 43.00 42.00 41.00 37.00 15.00 50.00 50.00 50.00 50.00 25.00 50.00 50.00 50.00 50.00 50.00
+F1% 44.10 43.04 42.33 39.18 16.95 46.00 44.59 43.54 39.32 16.77 31.17 30.84 30.71 27.94 13.94 32.17 31.38 31.16 28.14 14.31
+    The highest metrics are colored with gray. All metrics of UIED and Xianyu are lower than 0.01% and omitted.
 6.1.1. RQ1-1: Performance in terms of interactability
@@
 Consider the generally best model Orienter with Gemini, at the 0.75 threshold, the mAP peaks at 35.03%, indicating a strong ability to detect IGEs accurately.
-However, as the IoU threshold increases, indicating a stricter criterion for object detection, there is a noticeable decrease in Precision, dropping to 5.64% at an IoU of 0.95.
+However, as the IoU threshold increases, indicating a stricter criterion for object detection, there is a noticeable decrease in Precision, dropping to 5.64% at the IoU of 0.95.
 This trend suggests that while the Orienter is quite effective at a broader detection scope, its precision in highly specific object identification contexts is limited.
@@
 This trend of superiority is more significant at higher IoU thresholds.
-At the stringent IoU of 0.95, Orienter maintains a Precision of 19.02%, outperforming the closest baseline, CenterNet2 (Zhou et al., 2021), by over 321.7%, a significant margin considering the complexity of VR environments.
+At the stringent IoU of 0.95, Orienter maintains a Precision of 19.02%, outperforming the closest baseline, CenterNet2 (Zhou et al., 2021), by over 321.7%, a significant margin considering the complexity of XR environments.
 In terms of Recall and F1 Score, Orienter significantly outperforms other models by considerable margins.
 At an IoU of 0.75, Orienter’s Recall and F1 Score are 43% and 46.00%, respectively, surpassing CenterNet2 (Zhou et al., 2021)’s 30% and 29.32% by a notable margin of 43.3% and 56.9%, respectively.
-This indicates that Orienter is not only accurate but also reliable in identifying IGEs, a key requirement in industrial VR apps.
+This indicates that Orienter is not only accurate but also reliable in identifying IGEs, a key requirement in industrial XR apps.
 Note that the direct prompt method on LMMs performs poorly.
@@
 Similarly, at an IoU of 0.95, Orienter’s F1 Score of 14.31% far surpasses the closest baseline, YOLO v8 (Jocher et al., 2023), which scores 0.40%.
-This represents an increase of up to 3477.5%, highlighting Orienter’s ability to understand complex semantic structures within VR environments.
+This represents an increase of up to 3477.5%, highlighting Orienter’s ability to understand complex semantic structures within XR environments.
 The direct prompting method continues to perform poorly in terms of semantics due to the aforementioned challenges that LMM alone can not effectively address.
@@
 Our analysis demonstrates that Orienter significantly outperforms existing baseline models in both interactability and semantic understanding.
-The framework exhibits remarkable precision and recall rates, particularly in high IoU thresholds, which are critical for the nuanced and complex nature of industrial VR apps.
-This efficacy, especially in semantic understanding, underscores Orienter’s potential in revolutionizing IGE detection in VR environments.
+The framework exhibits remarkable precision and recall rates, particularly in high IoU thresholds, which are critical for the nuanced and complex nature of industrial XR apps.
+This efficacy, especially in semantic understanding, underscores Orienter’s potential in revolutionizing IGE detection in XR environments.
 6.1.3. RQ1-3: Performance in terms of context-sensitive interactability
+Table 3. Performance w.r.t. context-sensitive interactability, DL-based baselines are trained on our dataset*
+Faster-RCNN CenterNet2 Yolo v8 OmniParser gpt-4o-2024-08-06
+IoU P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1%
+0.75 2.44 0.14 0.27 4.27 0.27 0.51 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 2.44 0.07 0.13
+0.80 2.44 0.09 0.18 4.88 0.27 0.52 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 2.44 0.07 0.13
+0.85 ¡0.01 ¡0.01 ¡0.01 4.88 0.23 0.43 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+0.90 ¡0.01 ¡0.01 ¡0.01 4.88 0.18 0.34 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+0.95 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01 ¡0.01
+Gemini 1.5 Pro
+Orienter
+(gpt-4-vision-preview)
+Orienter
+(gpt-4o-2024-08-06)
+Orienter
+(Claude 3.5 Sonnet)
+Orienter
+(Gemini 1.5 Pro)
+IoU P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1%
+0.75 ¡0.01 ¡0.01 ¡0.01 44.72 6.66 10.77 67.70 24.00 32.51 62.74 18.54 26.07 67.91 22.96 31.40
+0.80 ¡0.01 ¡0.01 ¡0.01 44.57 6.32 10.41 67.99 23.36 31.82 62.93 17.94 25.36 67.91 22.20 30.63
+0.85 ¡0.01 ¡0.01 ¡0.01 45.02 6.27 10.40 67.82 22.21 30.62 64.17 16.84 24.27 68.42 21.40 29.75
+0.90 ¡0.01 ¡0.01 ¡0.01 45.24 5.37 9.26 68.93 19.35 27.75 63.96 14.84 21.84 70.01 18.23 26.78
+0.95 ¡0.01 ¡0.01 ¡0.01 26.83 1.62 3.03 47.56 8.99 13.59 47.07 7.47 11.37 53.35 7.01 11.76
 As shown in table 3, Orienter shows a promising performance to distinguish IGEs’ interactability in different contexts.
@@
 At the most stringent constraint of 0.95 IoU threshold, Orienter’s F1 score maintains over 10% while all metrics of baseline models are approximately 0, demonstrating Orienter’s significant advantage in understanding context-sensitive IGEs’ interactability against baseline models.
-Table 1. Performance of Orienter, DL-based baselines are trained on our dataset*
-Methods Interactability Semantics
-App Split Genre Split App Split Genre Split
-IoU 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95 0.75 0.80 0.85 0.90 0.95
-Faster- RCNN mAP18.05 13.84 7.32 1.47 0.07 17.90 12.88 6.96 1.56 0.13 0.21 0.15 0.03 0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
-P33.41 32.29 23.04 11.08 2.15 33.42 29.55 20.10 8.06 1.14 0.91 0.83 0.38 0.16 0.06 0.07 0.06 $\approx$0.0 $\approx$0.0 $\approx$0.0
-R32.00 25.00 18.00 8.00 1.00 29.00 22.00 15.00 6.00 1.00 4.00 4.00 3.00 3.00 4.00 2.00 2.00 $\approx$0.0 $\approx$0.0 $\approx$0.0
-F1% 32.69 28.18 20.21 9.29 1.36 31.05 25.22 17.18 6.88 1.06 1.48 1.37 0.68 0.30 0.11 0.13 0.12 $\approx$0.0 $\approx$0.0 $\approx$0.0
-Center- Net2 mAP17.79 14.98 11.73 5.91 0.93 15.63 12.53 9.07 4.23 0.45 0.20 0.18 0.15 0.05 $\approx$0.0 0.22 0.19 0.05 0.02 $\approx$0.0
-P32.53 28.52 26.85 18.26 4.83 28.67 25.81 27.71 16.46 4.51 0.68 0.65 0.57 0.27 0.05 0.62 0.59 0.27 0.19 0.04
-R32.00 29.00 23.00 15.00 3.00 30.00 27.00 18.00 12.00 3.00 4.00 4.00 3.00 3.00 2.00 5.00 5.00 3.00 3.00 3.00
-F1% 32.26 28.76 24.77 16.47 3.70 29.32 26.39 21.82 13.88 3.60 1.16 1.12 0.96 0.49 0.10 1.10 1.06 0.49 0.35 0.08
-Yolo v8 mAP15.30 13.12 9.85 6.16 1.51 11.54 10.02 7.59 3.86 0.46 0.02 0.02 0.01 $\approx$0.0 $\approx$0.0 0.03 0.02 0.02 0.01 0.01
-P26.58 25.38 20.63 16.65 7.31 23.92 21.38 24.22 14.55 2.68 0.21 0.20 0.19 0.05 0.01 0.29 0.28 0.27 0.25 0.21
-R26.00 23.00 20.00 14.00 3.00 23.00 22.00 15.00 11.00 3.00 4.00 4.00 4.00 1.00 3.00 3.00 3.00 3.00 3.00 3.00
-F1% 26.29 24.13 20.31 15.21 4.25 23.45 21.69 18.53 12.53 2.83 0.40 0.38 0.36 0.09 0.02 0.53 0.52 0.50 0.46 0.40
-[…]
-P0.21 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.43 0.43 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.25 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.23 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
-[…]
-F1% $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.47 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.46 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
-gpt-4o- 2024-08-06 mAP0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.01 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
-P0.78 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 0.75 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
-[…]
-F1% $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0
-Orienter (gpt-4-vision- preview) mAP25.53 24.32 23.03 19.64 4.61 27.07 25.22 23.87 19.80 4.54 14.93 14.66 13.96 12.17 3.67 15.96 15.57 14.62 12.87 3.75
-P34.79 33.52 32.32 29.31 14.92 38.42 36.70 34.97 31.40 16.79 16.06 16.05 15.17 15.27 6.53 17.11 16.85 15.78 15.93 6.14
-R40.00 39.00 38.00 34.00 12.00 40.00 38.00 37.00 32.00 11.00 50.00 50.00 50.00 33.00 20.00 50.00 50.00 50.00 33.00 20.00
-F1% 37.21 36.05 34.93 31.48 13.30 39.20 37.34 35.95 31.70 13.29 24.31 24.30 23.28 20.88 9.84 25.50 25.21 23.98 21.48 9.39
-Orienter (gpt-4o- 2024-08-06) mAP26.44 25.26 23.98 20.35 4.48 26.11 24.50 23.51 19.42 4.38 18.00 17.65 16.91 14.28 5.75 19.45 18.97 18.07 15.40 5.68
-P34.63 34.89 35.43 31.64 18.57 37.91 36.68 36.71 33.24 12.91 19.32 18.97 18.28 15.37 6.87 20.35 19.81 19.01 16.19 7.36
-R42.00 39.00 36.00 34.00 11.00 41.00 39.00 37.00 33.00 14.00 50.00 50.00 50.00 50.00 42.00 50.00 50.00 50.00 50.00 25.00
-F1% 37.96 36.83 35.71 32.78 13.81 39.39 37.80 36.85 33.12 13.43 27.87 27.50 26.77 23.52 11.81 28.93 28.38 27.55 24.46 11.37
-Orienter (Claude 3.5 Sonnet) mAP23.99 22.90 21.54 18.46 3.90 25.50 23.96 22.21 18.49 3.84 16.14 15.79 15.17 12.73 5.01 19.01 18.51 17.78 15.00 6.67
-P31.73 37.52 36.79 33.56 16.30 37.97 40.15 39.22 35.86 14.51 16.63 16.28 15.83 13.35 7.64 19.68 19.11 18.53 15.75 9.87
-R40.00 32.00 31.00 29.00 11.00 37.00 33.00 32.00 29.00 13.00 50.00 50.00 50.00 50.00 20.00 50.00 50.00 50.00 50.00 20.00
-F1% 35.39 34.54 33.65 31.11 13.14 37.48 36.23 35.24 32.07 13.71 24.95 24.56 24.04 21.07 11.06 28.24 27.66 27.04 23.96 13.22
-Orienter (Gemini 1.5 Pro) mAP32.71 31.27 30.51 26.33 5.88 35.03 32.59 30.82 25.08 5.64 20.69 20.35 20.19 17.77 7.16 22.12 21.44 21.17 18.45 7.41
-P46.42 45.30 44.94 41.64 21.46 49.45 47.52 46.42 41.94 19.02 22.64 22.29 22.16 19.39 9.66 23.72 22.86 22.64 19.58 8.35
-R42.00 41.00 40.00 37.00 14.00 43.00 42.00 41.00 37.00 15.00 50.00 50.00 50.00 50.00 25.00 50.00 50.00 50.00 50.00 50.00
-F1% 44.10 43.04 42.33 39.18 16.95 46.00 44.59 43.54 39.32 16.77 31.17 30.84 30.71 27.94 13.94 32.17 31.38 31.16 28.14 14.31
-    The highest metrics are colored with gray. Metrics below 0.01% are regarded as $\approx$ 0.0. All metrics of UIED and Xianyu are lower than 0.01% and omitted.
-Table 2. Performance w.r.t. context-sensitive interactability, DL-based baselines are trained on our dataset*
-Faster-RCNN CenterNet2 Yolo v8 gpt-4o-2024-08-06 Gemini 1.5 Pro
-Orienter
-(gpt-4-
-vision-preview)
-Orienter
-(gpt-4o-
-2024-08-06)
-Orienter
-(Claude 3.5
-Sonnet)
-Orienter
-(Gemini 1.5 Pro)
-IoU PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1%
-0.75 2.44 0.14 0.27 4.27 0.27 0.51 $\approx$0.0 $\approx$0.0 $\approx$0.0 2.44 0.07 0.13 $\approx$0.0 $\approx$0.0 $\approx$0.0 44.72 6.66 10.77 67.70 24.00 32.51 62.74 18.54 26.07 67.91 22.96 31.40
-0.80 2.44 0.09 0.18 4.88 0.27 0.52 $\approx$0.0 $\approx$0.0 $\approx$0.0 2.44 0.07 0.13 $\approx$0.0 $\approx$0.0 $\approx$0.0 44.57 6.32 10.41 67.99 23.36 31.82 62.93 17.94 25.36 67.91 22.20 30.63
-0.85 $\approx$0.0 $\approx$0.0 $\approx$0.0 4.88 0.23 0.43 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 45.02 6.27 10.40 67.82 22.21 30.62 64.17 16.84 24.27 68.42 21.40 29.75
-0.90 $\approx$0.0 $\approx$0.0 $\approx$0.0 4.88 0.18 0.34 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 45.24 5.37 9.26 68.93 19.35 27.75 63.96 14.84 21.84 70.01 18.23 26.78
-0.95 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 $\approx$0.0 26.83 1.62 3.03 47.56 8.99 13.59 47.07 7.47 11.37 53.35 7.01 11.76
+6.2. RQ2: Ablation Study
 Table 4. Contributions of different components*
@@
 Interactability Semantic Interactability Semantic Interactability Semantic Interactability Semantic
-IoU PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1% PRF1%
+IoU P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1% P% R% F1%
 0.75 47.14 43.00 44.97 21.36 50.00 29.93 38.27 39.00 38.63 18.14 50.00 26.62 38.51 36.00 37.21 18.85 50.00 27.38 31.40 29.00 30.15 16.78 50.00 25.13
@@
 0.95 19.25 16.00 17.47 8.18 50.00 14.06 13.80 14.00 13.90 6.86 33.00 11.36 13.91 13.00 13.44 6.82 33.00 11.30 11.58 10.00 10.73 7.22 16.00 9.95
-6.2. RQ2: Ablation Study
-Table 3 shows the performance of Orienter and its three variations.
-For Orienter w/o Context Comprehension, we remove the whole Module I Semantic Context Comprehension.
-For Orienter w/o Reflection-Directed Loop, we remove the whole Module II instead of IGE candidate detection module (to at least localize elements).
-For Orienter w/o interactability Classification, we remove the whole Module III Context-Sensitive Interactability Classification.
+Table 4 shows the performance of Orienter and its three variations. We remove the corresponding components of each variation to demonstrate their neceesity.
 As can be seen in Table 4, Orienter outperforms all its variants significantly, demonstrating the necessity of the designed pipeline.
@@
 The experimental results show the necessity of each component in Orienter, highlighting their critical role in accurately detecting IGEs.
-6.3. RQ3: Usefulness
+6.3. RQ3: Effectiveness for boosting automated testing
+Table 5. Average testing result by apps with different categories. Categories with # of app ¡ 10 are omitted
+App Category # App # GUI Scene Random IGE Cov. Guided IGE Cov. Improved
+All 77 936 0.65 0.83 +26.91%
+Singleplayer 43 640 0.45 0.60 +32.42%
+Casual 46 600 0.42 0.52 +24.75%
+Simulation 30 521 0.37 0.49 +31.05%
+3D 24 471 0.35 0.47 +33.43%
+Immersive Sim 13 374 0.28 0.38 +35.05%
+First-Person 28 402 0.29 0.36 +25.61%
+Puzzle 14 339 0.25 0.34 +36.60%
+Free to Play 36 393 0.27 0.33 +20.64%
+Early Access 11 319 0.24 0.32 +32.52%
+Indie 38 407 0.27 0.32 +19.88%
+Colorful 13 318 0.24 0.32 +35.39%
+6DOF 14 295 0.21 0.27 +29.22%
+Education 12 292 0.21 0.27 +31.40%
+Atmospheric 24 267 0.18 0.21 +15.05%
+Adventure 23 224 0.15 0.19 +26.29%
+Action 24 220 0.15 0.17 +13.06%
+Physics 14 178 0.12 0.15 +21.87%
+Exploration 14 181 0.12 0.15 +21.17%
+Arcade 17 163 0.11 0.13 +25.59%
+Funny 10 93 0.05 0.07 +34.46%
+Strategy 11 87 0.06 0.07 +13.52%
+Sci-fi 11 65 0.04 0.04 +8.84%
+Multiplayer 11 56 0.03 0.03 -11.37%
+Mean 20 300 0.21 0.27 +24.19%
 (a) Effective interaction count
 (b) Coverage rate
-Figure 14. Results of RQ3: Usefulness
+Figure 15. Performance of testing over time across strategies
 As shown in Figure 15, the testing guided by Orienter outperforms the random testing with a significant gap.
@@
 The number of effective interactions in guided testing increases notably faster than in random generation.
-At the end of the testing, the guided one’s average total number of effective interactions reaches 6,571.8, surpassing the random testing’s 2,911.6 by a remarkable 125.7%.
-Regarding the rate of effective interaction, guided testing peaks at 0.57, with an average of 0.46 over the whole duration of testing, considerably better than that of the random testing remaining at around 0.11.
-At 10 minutes, the IGE coverage of the guided testing already reaches 0.65, outperforming the random testing’s 0.32 by 103.1%.
-The IGE coverage of the guided testing eventually reaches 0.83, which is 27.7% higher than the random testing’s 0.65.
-These results demonstrate the capability of Orienter to boost the effectiveness and efficiency of VR app testing.
-Apart from automated VR app testing, other downstream software engineering problems like UI to code, video-based or screenshot-based testing, etc., can also benefit from the ability of Orienter to effective and efficient IGE detection.
+At the end of the testing, the guided one’s average number of effective interactions reaches 13,038.03 on 936 GUI Scenes in total, surpassing the random testing’s 5,911.53 by a remarkable 120.6%.
+At 10 minutes, the IGE coverage of the guided testing already reaches 0.64, surpassing the random testing’s 0.30 by 113.3%.
+The guided testing eventually reaches a coverage of 0.83, 27.7% higher than the random testing’s 0.65.
+Figure 16. Coverage rate on apps in different categories. Categories with # of app ¡ 10 are omitted
+We further investigate the performance of guided and non-guided testing agents on apps with different categories.
+We select 23 community-generated categories with no less than 10 apps (except for the category ”VR”) for category-wise evaluation in terms of IGE coverage, covering 76 of the 77 AUTs.
+Table 5 and Figure 16 intuitively reveal the benefits of the guidance from Orienter.
+The testing agent with the guidance of Orienter surpasses the one without guidance in most app categorys, achieving improvements by up to 36.60% (0.34 vs.
+0.25) in the category ”Puzzle”.
+The guided agent achieves the highest IGE coverage of 0.60 in the category ”Singleplayer”, surpassing the random agent by 32.42%.
+In contraction, the guided testing agent shows a worse yet still very close IGE coverage compared with the random testing agent in the ”Multiplayer” category, both achieving a coverage at around 0.03.
+Such reduction suggests that the bias introduced by the external guidance could also lead to performance degradation for testing agents, highlighting the necessity of ”thinking outside the box” to explore the GUI scene aside from the guidance to some extent.
+The large gap between the two opposite categories also reflects the huge difference between different categories of apps.
+In addition, the guided testing agent generally performs more consistently than the random testing agent between testing rounds.
+These results demonstrate the capability of Orienter to boost the effectiveness and efficiency of XR app testing.
+Apart from automated XR app testing, other downstream software engineering problems like UI to code, video-based or screenshot-based testing, etc., can also benefit from the ability of Orienter to effective and efficient IGE detection.
 7. Threats to Validity
@@
 An internal threat to validity is related to the annotators’ bias and subjective interpretation.
-Given that the dataset was annotated by 13 annotators, there’s potential bias based on their personal experience and interpretation of interactability in VR apps.
+Given that the dataset was annotated by 13 annotators, there’s potential bias based on their personal experience and interpretation of interactability in XR apps.
 This could impact the consistency and accuracy of the annotations.
@@
 External Validity.
-One threat to external validity is the generalizability across VR apps.
-The diversity in VR apps poses a challenge to generalizability.
-Orienter’s effectiveness demonstrated in specific scenarios may not uniformly apply across different VR environments, particularly those with unique interaction paradigms or novel GUI element types.
-To mitigate this threat, we sample popular and complicated VR apps, crossing a large portion of app categories.
+One threat to external validity is the generalizability across XR apps.
+The diversity in XR apps poses a challenge to generalizability.
+Orienter’s effectiveness demonstrated in specific scenarios may not uniformly apply across different XR environments, particularly those with unique interaction paradigms or novel GUI element types.
+To mitigate this threat, we sample popular and complicated XR apps, crossing a large portion of app categories.
 8. Related Work
-8.1. Studies on VR/AR Apps
+8.1. IGE Detection
+Empirical studies have been conducted to give insights into IGE detection (Ye et al., 2021), e.g., Chen et al.
+(Chen et al., 2020) conducted a systematic large-scale empirical study on GUI element detection methods.
+Previous works have adapted both old-fashioned methods  (Yeh et al., 2009; Bao et al., 2015; Dixon and Fogarty, 2010) and deep learning models (Wang et al., 2023a; Qian et al., 2022; Wu et al., 2023) to IGE detection.
+White et al.
+(White et al., 2019) replaced methods that involve GUI APIs with machine learning techniques to find interactable elements in GUI images.
+Xie et al.
+(Xie et al., 2020) proposed UIED, which combines old-fashioned computer vision approaches and deep learning models to detect components on complex GUI images.
+For game apps, Ye et al.
+(Ye et al., 2021) conducted an empirical study of clickable GUI element detection on mobile games and constructed the first clickable GUI element detection benchmark.
+Wu et al.
+(Wu et al., 2023) constructed a larger and more precise GUI dataset for mobile games and evaluated seven IGE detection techniques on it.
+The industry also applies IGE detection techniques to solve business problems like GUI testing (Ran et al., 2022; Wang et al., 2023a; Qian et al., 2022).
+Ran et al.
+(Ran et al., 2022) reported their experiences and lessons learned developing and deploying VTest, an automated visual testing framework for smartphones that leverages only GUI images.
+Wang et al.
+(Wang et al., 2023a) proposed iExplorerGame, a unified framework for black-box mini-game testing that combines deep learning object detection and edge aggregation-based segmentation to guide the testing.
+Qian et al.
+(Qian et al., 2022) proposed a fast OCR-based widget localization technique Label Text Screening, which accelerates the OCR widget detection by analyzing and leveraging the feature of texts in widgets.
+Research on UI analysis is going in-depth and reaching fields with complex GUI-like game apps.
+However, existing methods are limited to common IGE with finite predefined categories, which fail to tackle the open vocabulary challenge.
+8.2. VR/AR Testing
+VR/AR (XR) apps provide an immersive experience to users, involving various procedures like device tracking and rendering.
+To ensure users’ experience, there are several works conducting empirical studies on XR software testing (Andrade et al., 2019).
+Andrade et al.
+(Andrade et al., 2019) compared open-source VR projects with non-VR projects to point out the necessity of performing testing on VR apps.
+Rzig et al.
+(Rzig et al., 2023) performed a large-scale empirical study on software testing practices of open-source VR projects.
+Several works have proposed approaches to facilitate XR testing.
+Qin and Hassan (Qin and Hassan, 2022) proposed DyTRec, which provides developers with recommendations on which codes should be tested by actually running the app.
+Souza et al.
+(Souza et al., 2018) proposed VR-ReST, a tool designed to assist with requirements specification and test data generation for VR apps.
+Rafi et al.
+(Rafi et al., 2022) proposed PredART which can be used as the test oracle when checking the placement of virtual objects in AR.
+Wang (Wang, 2022) proposed VRTest that extracts information from the VR scene and automatically explores the scene and interacts with the objects by controlling the camera.
+Andrade et al.
+(de Andrade et al., 2023) proposed a novel approach that combines metamorphic testing, agent-based testing, and machine learning to test VR apps.
+Recently, Li et al.
+(Li et al., 2024) proposed StereoID to automatically detect stereoscopic visual inconsistencies in VR apps.
+8.3. Exploratory Studies on VR/AR Apps
 Previous works have conducted several empirical studies to understand VR/AR (XR) apps from different perspectives.
@@
 (Guo et al., 2024) developed a security and privacy assessment tool and conducted an empirical study on Oculus VR apps.
-8.2. VR/AR Testing
-VR/AR (XR) apps provide an immersive experience to users, involving various procedures like device tracking and rendering.
-To ensure users’ experience, there are several works conducting empirical studies on XR software testing (Andrade et al., 2019).
-Andrade et al.
-(Andrade et al., 2019) compared open-source VR projects with non-VR projects to point out the necessity of performing testing on VR apps.
-Rzig et al.
-(Rzig et al., 2023) performed a large-scale empirical study on software testing practices of open-source VR projects to identify the current state of testing in VR apps.
-Several works have proposed approaches to facilitate XR testing.
-Qin and Hassan (Qin and Hassan, 2022) proposed DyTRec, which provides developers with recommendations on which codes should be tested by actually running the app.
-Souza et al.
-(Souza et al., 2018) proposed VR-ReST, a tool designed to assist with requirements specification and test data generation for VR apps.
-Rafi et al.
-(Rafi et al., 2022) proposed PredART that can be used as the test oracle when checking the placement of virtual objects in AR.
-Wang (Wang, 2022) proposed VRTest that extracts information from the VR scene and automatically explores the scene and interacts with the objects by controlling the camera.
-Andrade et al.
-(de Andrade et al., 2023) proposed a novel approach that combines metamorphic testing, agent-based testing, and machine learning to test VR apps.
-Recently, Li et al.
-(Li et al., 2024) proposed StereoID to automatically detect stereoscopic visual inconsistencies in VR apps.
-8.3. User Interface Analysis
-Empirical studies have been conducted to give insights into UI analysis (Ye et al., 2021), such as Chen et al.
-(Chen et al., 2020) conducted a systematic large-scale empirical study on GUI element detection methods.
-Previous works have adapted both old-fashioned methods  (Yeh et al., 2009; Bao et al., 2015; Dixon and Fogarty, 2010) and deep learning models (Wang et al., 2023a; Qian et al., 2022; Wu et al., 2023) to UI analysis.
-White et al.
-(White et al., 2019) replaced methods that involve GUI APIs with machine learning techniques to find interactable elements in GUI images.
-Xie et al.
-(Xie et al., 2020) proposed UIED, which combines old-fashioned computer vision approaches and deep learning models to detect components on complex GUI images.
-For game apps, Ye et al.
-(Ye et al., 2021) conducted an empirical study of clickable GUI element detection on mobile games and constructed the first clickable GUI element detection benchmark.
-Wu et al.
-(Wu et al., 2023) constructed a larger and more precise GUI dataset for mobile games and evaluated seven GUI detection techniques on it.
-The industry also applies UI analysis techniques to solve business problems like GUI testing.
-Ran et al.
-(Ran et al., 2022) reported their experiences and lessons learned developing and deploying VTest, an automated visual testing framework for smartphones that leverages only GUI images.
-Wang et al.
-(Wang et al., 2023a) proposed iExplorerGame, a unified framework for black-box mini-game testing that combines deep learning object detection and edge aggregation-based segmentation to guide the testing.
-Qian et al.
-(Qian et al., 2022) proposed a fast OCR-based widget localization technique Label Text Screening, which accelerates the OCR widget detection by analyzing and leveraging the feature of texts in widgets.
-Research on UI analysis is going in-depth and reaching fields with complex GUI-like game apps.
-However, existing methods are limited to common IGE with finite predefined categories, which fail to tackle the open vocabulary challenge, as mentioned in  section 1.
 9. Conclusion
-In this paper, we analyze the key challenges in interactable GUI element (IGE) detection for Virtual Reality (VR) apps.
-We propose Orienter, the first zero-shot, context-sensitive IGE detection framework for VR apps.
+In this paper, we analyze the key challenges in interactable GUI element (IGE) detection for Extended Reality (XR) apps.
+We propose Orienter, the first zero-shot, context-sensitive IGE detection framework for XR apps.
 Extensive experiments have been conducted to verify the effectiveness of the proposed IGE detection framework.
 The results demonstrate that Orienter is more effective than the state-of-the-art GUI element detection approaches, Experiments also illustrate that Orienter is beneficial for boosting automated GUI testing.
-Alongside developing Orienter, we have also created the first dedicated dataset for IGE detection in VR environments, both of which are publicly available to spur further research in this field.
+Alongside developing Orienter, we have also created the first dedicated dataset for IGE detection in XR environments, both of which are publicly available to spur further research in this field.